### v1

# references.bib

Abstract
We study the convergence of gradient methods for the training of mean-field single hidden layer neural networks with square loss. Observing this is a separable non-linear least-square problem which is linear w.r.t. the outer layer’s weights, we consider a Variable Projection (VarPro) or two-timescale learning algorithm, thereby eliminating the linear variables and reducing the learning problem to the training of the feature distribution. Whereas most convergence rates for the training of neural networks rely on a neural tangent kernel analysis where features are fixed, we show such a strategy enables provable convergence rates for the sampling of a teacher feature distribution. Precisely, in the limit where the regularization strength vanishes, we show that the dynamic of the feature distribution corresponds to a weighted ultra-fast diffusion equation. Relying on recent results on the asymptotic behavior of such PDEs, we obtain guarantees for the convergence of the trained feature distribution towards the teacher feature distribution in a teacher-student setup.
1 Introduction
Machine learning methods based on artificial neural networks have recently experienced a significant increase in popularity due to their efficiency in solving numerous supervised or unsupervised learning tasks. This success owes to their capacity to perform feature learning, that is to extract meaningful representations from the data during the training process [Goodfellow-et-al-2016, Chap. 15], standing in contrast with kernel methods for which feature representations are designed by hand and fixed during training [hofmann2008kernel]. Feature learning is believed to play a fundamental role in the generalization performance of neural networks. For example, adaptivity to low-dimensional representations of the data can prevent the curse of dimensionality [bach2017breaking].
However, the process through which features are learned remains largely misunderstood. Indeed, adaptability of the representation comes in neural networks at the price of a nonlinear parameterization, making the training dynamic more difficult to analyze. Specifically, for overparameterized neural networks architecture where the dimension of the parameter space greatly exceeds the number of training samples, recent works have put forward the crucial role played by the choice of scaling w.r.t. the number of parameters in the training dynamic [chizat2019lazy, liu2020linearity, yang2021tensor]. For single hidden layer neural networks, the “kernel regime”, corresponding to a scaling of $1/\sqrt{M}$ where $M$ is the width, has been identified as a scaling for which the model is well-approximated by its linearization around initialization, therefore reducing to a kernel method [jacot_neural_2021]. Relying on the good conditioning of the “Neural Tangent Kernel (NTK)”, this regime provides convergence of gradient descent towards a global minimizer of the risk at a linear rate [allen2019convergence, du2019gradient, lee2019wide, zou_gradient_2020]. However, this regime has also been shown to suffer from a “lazy training” behavior preventing significant modification of the feature distribution and associated to poor generalization guarantees [chizat2019lazy].
In contrast, an other line of work has been focused on the “mean field” regime corresponding to a scaling of $1/M$ (which we consider in eq. 1) for which the neural network is parameterized by a probability distribution over the space of weights [chizat2018global, mei2019mean, rotskoff2019global, sirignano2020mean]. While such a choice of scaling has been shown to enable nonlinear feature learning behaviors [yang2021tensor], existing convergence results are primarily qualitative, lacking explicit convergence rates. To bridge this gap, we are interested in this work in the dynamic of the feature distribution in the training of mean-field models of shallow neural network architectures. We study more particularly a variable projection or two-timescale learning strategy which allows reducing the learning problem to the training of the feature distribution.
1.1 Mean-field neural networks and two-timescale learning
We consider in this work neural networks with parameter in the parameter space $\Omega$, which is assumed to be either the $n$-dimensional torus $\mathbb{T}^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$, or a closed, bounded and convex domain of $\mathbb{R}^{n}$. In the following, we are given a map $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$, called feature map, and for an integer $M\geq 1$, we define a Single Hidden Layer (SHL) neural network of width $M$ with inner weights $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ and outer weights $\{u_{i}\}_{1\leq i\leq M}\in\mathbb{R}^{M}$ as the map:
$F_{\{(\omega_{i},u_{i})\}}:x\in\mathbb{R}^{d}\mapsto\frac{1}{M}\sum_{i=1}^{M}u_{i}\phi(\omega_{i},x)\in\mathbb{R},$ (1)
taking inputs in the input space $\mathbb{R}^{d}$ and returning values in the output space $\mathbb{R}$. Thanks to the interchangeability of the indices and to the normalisation factor $1/M$, the above model can be reparameterized in terms of the empirical distribution of the inner weights $\{\omega_{i}\}_{1\leq i\leq M}$. Given an arbitrary probability distribution $\mu\in\mathcal{P}(\Omega)$ on the space of inner weights and a measurable map $u\in L^{1}(\mu)$ we define:
$F_{\mu,u}:x\in\mathbb{R}^{d}\mapsto\int_{\Omega}u(\omega)\phi(\omega,x)\mathrm{d}\mu(\omega)\in\mathbb{R}.$ (2)
In particular, for the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$ and the outer weights $\hat{u}(\omega_{i})=u_{i}$ we recover the finite width SHL $F_{\hat{\mu},\hat{u}}=F_{\{(\omega_{i},u_{i})\}}$. Such a “mean-field” model of neural network has been proposed by several authors to study the training of neural network at arbitrary large width [chizat2018global, mei2019mean, rotskoff2019global, sirignano2020mean].
We consider in this work a regression setting where the neural network weights are trained for minimizing the mean square error between the neural network’s output and a target signal. In the following, we are given an input data distribution $\rho\in\mathcal{P}(\mathbb{R}^{d})$, a target signal $Y\in L^{2}(\rho)$ and we define the training risk for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$ as:
$\mathcal{R}(\mu,u)\coloneqq\frac{1}{2}\|F_{\mu,u}-Y\|^{2}_{L^{2}(\rho)}\,.$
Training the neural networks amounts to find parameters $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$ minimizing the training risk $\mathcal{R}(\mu,u)$. Such a task is usually performed using first order optimization methods such gradient descent or stochastic versions of gradient descent on the neural network’s weights [bottou2018optimization].
Variable Projection
We study a Variable Projection (VarPro) algorithm where elimination of the linear variable $u$ enables reducing the training of a neural network to the sole problem of learning the feature distribution. Introduced in [golub1973differentiation] for the minimization of separable nonlinear least-squares problems, such a strategy has proven to be efficient in various applications [golub2003separable, osborne2007separable]. The principle of this method can however apply to a broader class of problems [osborne2007separable, aravkin2012estimating] and, for general convex functions, one can show using Schur complement that partial optimization over one variable leads to a better conditioning of the Hessian [calafiore2014optimization, vialard2022partial].
Exploiting the linearity w.r.t. the outer weights in the definition of $F$, it is convenient to read a neural network’s output $F_{\{(\omega_{i},u_{i})\}}(x)=\frac{1}{M}\sum u_{i}\phi(\omega_{i},x)$ as a linear combination of the features $\{\phi(\omega_{i},x)\}_{i=1}^{M}$. From this point of view, neural networks should be compared to kernel methods for which the features are built in advance and fixed during training, whereas only the weights of the linear combination are learned [hofmann2008kernel]. In contrast, both inner weights $\{\omega_{i}\}_{i=1}^{M}$ and outer weights $\{u_{i}\}_{i=1}^{M}$ of a neural network are usually trained. In the following, we refer to the parameters $\omega\in\Omega$ as the neural network’s features and to $\mu\in\mathcal{P}(\Omega)$ as the feature distribution. More generally in the mean-fied limit, for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$, we have:
$F_{\mu,u}=\int_{\Omega}\phi(\omega,.)u(\omega)\mathrm{d}\mu(\omega)=\Phi_{\mu}\cdot u$ (3)
where we introduced the feature operator $\Phi_{\mu}:u\in L^{1}(\mu)\mapsto\int_{\Omega}u(\omega)\phi(\omega,.)\mathrm{d}\mu(\omega)\in L^{2}(\rho)$. Given a regularization parameter $\lambda>0$, we consider the regularization $\mathcal{R}^{\lambda}$ of the training risk w.r.t. the outer weights defined for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$ as:
$\mathcal{R}^{\lambda}(\mu,u)\coloneqq\frac{1}{2}\|F_{\mu,u}-Y\|^{2}_{L^{2}(\rho)}+\lambda\|u\|^{2}_{L^{2}(\mu)}=\frac{1}{2}\|\Phi_{\mu}\cdot u-Y\|^{2}_{L^{2}(\rho)}+\lambda\|u\|^{2}_{L^{2}(\mu)},$
where we assume $\mathcal{R}^{\lambda}(\mu,u)=+\infty$ if $u\notin L^{2}(\mu)$. One can notice that the problem of minimizing the risk $\mathcal{R}^{\lambda}$ belongs to the class of separable nonlinear least-squares problems. Indeed, given a fixed inner weights’ distribution $\mu\in\mathcal{P}(\Omega)$, the problem of minimizing $\mathcal{R}^{\lambda}$ w.r.t. $u$ is a ridge regression problem and, for $\lambda>0$, there exists a unique solution $u^{\lambda}[\mu]\in\operatorname*{arg\,min}_{u\in L^{2}(\mu)}\mathcal{R}^{\lambda}(\mu,u)$ given by $u^{\lambda}[\mu]\coloneqq(\Phi_{\mu}^{\top}\Phi_{\mu}+2\lambda)^{-1}\Phi_{\mu}^{\top}Y$. Plugging this in $\mathcal{R}^{\lambda}$ gives rise to a reduced risk defined for any $\mu\in\mathcal{P}(\Omega)$ by:
$\mathcal{L}^{\lambda}(\mu)\coloneqq\frac{1}{\lambda}\mathcal{R}^{\lambda}(\mu,u^{\lambda}[\mu])=\min_{u\in L^{2}(\mu)}\frac{1}{2\lambda}\|\Phi_{\mu}\cdot u-Y\|^{2}_{L^{2}(\rho)}+\|u\|^{2}_{L^{2}(\mu)}.$ (4)
This definition also extends to the limiting case $\lambda\to 0^{+}$ by considering:
$\mathcal{L}^{0}(\mu)\coloneqq\min_{\Phi_{\mu}\cdot u=Y}\|u\|^{2}_{L^{2}(\mu)}.$ (5)
where the infimum is taken to be +$\infty$+ ∞ whenever the signal $Y$ is not in the range of $\Phi_{\mu}$. In the case where $Y\in\operatorname{Range}(\Phi_{\mu})$, this minimization problem admits a unique solution and $\mathcal{L}^{0}(\mu)=\|u^{\dagger}[\mu]\|^{2}_{L^{2}(\mu)}$, where $u^{\dagger}[\mu]=\Phi_{\mu}^{\dagger}Y$ and $\Phi_{\mu}^{\dagger}$ is the generalized pseudo-inverse of $\Phi_{\mu}$ restricted to $L^{2}(\mu)$.
Two-timescale learning
We study the behavior of first order methods such as gradient flow or gradient descent for the minimization of the reduced risk $\mathcal{L}^{\lambda}$ for $\lambda\geq 0$. Whereas the algorithmic complexity of gradient descent might be high as the evaluation of $\mathcal{L}^{\lambda}$ implies the resolution of a linear regression problem, it can also conveniently be be seen as a two-timescale limit of gradient descent performed on the risk $\mathcal{R}^{\lambda}$ [borkar1997stochastic, borkar2008stochastic].
For a neural network of finite width $M\geq 0$ with features $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$, the associated reduced risk is ^$\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}\}_{1\leq i\leq M})\coloneqq\mathcal{L}^{\lambda}(\hat{\mu})$ where $\hat{\mu}$ is the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. Then, for an initialization $\{\omega_{i}^{0}\}_{1\leq i\leq M}\in\Omega^{M}$ and a step-size $\eta>0$, the gradient descent algorithm updates the features iteratively following:
$\forall k\geq 0,\forall i\in\{1,...,M\},\quad\omega_{i}^{k+1}=\omega_{i}^{k}-\eta\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}^{k}\}_{1\leq i\leq M})\,.$
For the purpose of theoretical analysis we consider the limit of the gradient descent algorithm when the step-size $\eta$ tends to 00. For an initialization $\{\omega_{i}(0)\}_{1\leq i\leq M}\in\Omega^{M}$ the gradient flow of ^$\hat{\mathcal{L}}^{\lambda}_{f}$ is defined for every time $t\geq 0$ as the solution $\{\omega_{i}(t)\}_{1\leq i\leq M}\in\Omega^{M}$ to the ODE:
$\forall i\in\{1,...,M\},\quad\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)=-\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}(t)\}_{1\leq i\leq M})\,.$ (6)
Similarly, for a neural network with weights $\{(\omega_{i},u_{i})\}_{1\leq i\leq M}\in(\Omega\times\mathbb{R})^{M}$ the associated risk is ^$\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i},u_{i})\}_{1\leq i\leq M})\coloneqq\mathcal{R}^{\lambda}(\hat{\mu},\hat{u})$, where $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$ and $\hat{u}(\omega_{i})=u_{i}$, and for an initialization $\{(\omega_{i}(0),u_{i}(0))\}_{1\leq i\leq M}$, the gradient flow dynamic reads:
$\forall i\in\{1,...,M\},\quad\left\{\begin{array}[]{cc}\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)&=-\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})\\[8.0pt] \frac{\mathrm{d}}{\mathrm{d}t}u_{i}(t)&=-\gamma\nabla_{u_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})\end{array}\right.$
where we introduce the timescale parameter $\gamma>0$ controlling the ratio of learning timescales between the parameters $\{\omega_{i}\}_{1\leq i\leq M}$ and $\{u_{i}\}_{1\leq i\leq M}$. When $\gamma<1$ the outer-weights $u_{i}$ are learned “slower” than the features $\omega_{i}$ and reversely, when $\gamma>1$ the outer-weights $u_{i}$ are learned “faster” than the features $\omega_{i}$. In particular, the limiting training dynamic when $\gamma\to+\infty$ corresponds (formally) to the case where the outer weights are learned “instantaneously”, that is, at each time $t\geq 0$, we have $\{u_{i}(t)\}_{1\leq i\leq M}\in\operatorname*{arg\,min}_{u\in\mathbb{R}^{M}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i})\}_{1\leq i\leq M})$. Then by the envelope theorem $\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})=\lambda\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}(t)\}_{1\leq i\leq M})$. The training dynamic eq. 6 can thus be obtained (at least formally) as the training dynamic of a neural network’s features in the limit where $\gamma\to+\infty$. We refer to eq. 6 as the two-timescale regime.
Wasserstein gradient flows and ultra-fast diffusions
Relying on the mathematical framework provided by theory of gradient flows in the space of probability measures [ambrosio2008gradient, santambrogio2017euclidean], we show in section 4 that the dynamic of the feature distribution when trained with gradient flow for the minimization of the reduced risk $\mathcal{L}^{\lambda}$ is solution to an advection PDE of the form:
$\partial\mu_{t}-\mathrm{div}(\mu_{t}\nabla\mathcal{L}^{\lambda}[\mu_{t}])=0$
for some nonlinear velocity field $\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]$. We study in section 5 the asymptotics of this equation when the training time $t$ tends to +$\infty$+ ∞ and the regularization strength $\lambda$ tends to $0^{+}$. We are more particularly interested in the case where the signal $Y$ itself can be exactly represented by a neural network. We consider the following assumption:
Assumption 1 (Teacher student setup).
We assume that,
  1. 1.
    there exists a measure $\bar{\nu}\in\mathcal{M}(\Omega)$ s.t. $Y=\Phi\star\bar{\nu}$, where $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ is the convolution operator defined by:
    $\forall\nu\in\mathcal{M}(\Omega),\quad\Phi\star\nu\coloneqq\int_{\Omega}\phi(\omega,.)\mathrm{d}\nu(\omega)\;;$ (7)
  2. 2.
    the operator $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ is injective.
In this case, we refer to $\bar{\nu}\in\mathcal{M}(\Omega)$ as the teacher measure and to $\bar{\mu}\coloneqq|\bar{\nu}|/\|\bar{\nu}\|_{\mathrm{TV}}\in\mathcal{P}(\Omega)$ as the teacher (feature) distribution.
In such a “teacher-student” framework, we are interested in determining to what extent the teacher feature distribution can be learned by the student neural network. Observe that, under Assumption 1, $\mathcal{L}^{0}$ can be simply expressed in terms of the $\chi^{2}$-divergence between the teacher feature distribution $\bar{\mu}$ and $\mu$. By definition $\chi^{2}(\bar{\mu}|\mu)=\int_{\Omega}\left|\frac{\mathrm{d}\bar{\mu}}{\mathrm{d}\mu}-1\right|^{2}\mathrm{d}\mu$ and:
$\mathcal{L}^{0}(\mu)=\int_{\Omega}\left|\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu}\right|^{2}\mathrm{d}\mu=\|\bar{\nu}\|^{2}_{\mathrm{TV}}\left(\int_{\Omega}\left|\frac{\mathrm{d}\bar{\mu}}{\mathrm{d}\mu}-1\right|^{2}\mathrm{d}\mu+1\right)=\|\bar{\nu}\|^{2}_{\mathrm{TV}}(\chi^{2}(\bar{\mu}|\mu)+1)\,.$
The Wasserstein gradient flow of $\mathcal{L}^{0}$ then corresponds to nonlinear diffusion equations of the form:
$\partial_{t}\mu=\mathrm{div}\left(\bar{\mu}\nabla\left(\frac{\mu}{\bar{\mu}}\right)^{m}\right)$ (8)
[…]
Remark 1.1.
Some remarks about Assumption 1:
  • At fixed $\lambda>0$, the teacher-student assumption that $Y=\Phi\star\bar{\nu}$ is not restrictive since one can always replace $Y$ by its orthogonal projection on the set $\{\Phi\star\nu,\,\nu\in\mathcal{M}(\Omega)\}$, thereby only modifying $\mathcal{L}^{\lambda}$ by subtracting a constant term. However, this assumption becomes crucial in the limit $\lambda\to 0^{+}$ to ensure the feasibility of the the optimization problem in eq. 10.
  • The injectivity assumption on $\Phi\star$ ensures uniqueness of the reference measure $\bar{\nu}$. In the limit where $\lambda\to 0^{+}$, this allows to rewrite the $\mathcal{L}^{0}$ only in terms of a divergence between $\bar{\nu}$ and $\mu$ (eq. 17). In the case $\lambda>0$, $\mathcal{L}^{\lambda}$ is an infimal convolution between this divergence and a kernel discrepency (eq. 18) and the injectivity assumption ensures this discrepency is a distance on the space of measures (Lemma A.1). It will be useful in section 5 to prove convergence of Wasserstein gradient flows of $\mathcal{L}^{\lambda}$ to solutions of ultra-fast diffusion equations. In the case of a two-layer neural network when the feature map is of the form $\phi((w,b),x)=\sigma(w^{\top}x+b)$ the injectivity assumption is satisfied as soon as $\sigma$ is not a polynomial and the data distribution has full support on $\mathbb{R}^{d}$ ([sun2019random, Thm. III.4]).
1.2 Contributions and related works
Contributions
This paper studies the convergence of gradient methods for the training of a mean-field model of neural network in the two-timescale regime. Precisely, we study the dynamic of the feature distribution $\mu\in\mathcal{P}(\Omega)$ when trained with gradient flow for the minimization of the reduced risk $\mathcal{L}^{\lambda}$, for $\lambda\geq 0$. In the teacher-student scenario defined by Assumption 1, we establish guarantees for the convergence of $\mu$ towards the teacher feature distribution $\bar{\mu}$:
  • In the case $\lambda=0$, we show in section 4 that this dynamic corresponds to an ultra-fast diffusion equation. Relying on the work of [iacobelli2019weighted], this allows to state convergence towards the teacher feature distribution $\bar{\mu}$ (Theorem 4.3), with a linear convergence rate.
  • At fixed $\lambda>0$, we establish in Theorem 5.1 convergence of $\mu$ towards the teacher feature distribution $\bar{\mu}$ with an algebraic rate.
  • Finally, in the limit $\lambda\to 0^{+}$, we show under regularity assumptions that, in the two-timescale regime, the dynamic of the feature distribution $\mu$ converges locally uniformly in time to the solution of the ultra-fast diffusion equation with weights $\bar{\mu}$ (Theorem 5.2).
Convergence analysis for the training mean-field neural networks
Several works have studied the convergence of gradient based methods for the training of neural networks models similar to eq. 1 with the mean-field scaling $\frac{1}{M}$. In [bach2021gradient], the authors show that for two layer neural networks with a homogeneous activation, if gradient flow on the weights distribution converges then it converges towards a global minimizer of the risk. In [rotskoff2019global], a similar result is shown for a modification of the gradient flow dynamic where a supplementary “birth-death” term is added.
Several works have also analyzed the convergence of noisy gradient descent, or Langevin dynamic, for the training of mean-field models of two layer neural networks [chizat2022mean, mei2019mean, nitanda2022convex, hu2021mean, suzuki2023feature]. Thanks to the addition of an entropic regularization, these work provide a convergence rate for the sampling of an invariant weight distribution.
Two-timescale learning
While two-timescales learning strategies have a broad range of application in the fields of stochastic approximation and optimization [borkar1997stochastic, borkar2008stochastic], there has been a recent interest in these methods for the training of neural networks [marion2023leveraging, berthier2024learning, wang2024mean, bietti2023learning, takakura2024mean]. Specifically, [berthier2024learning] studies the training of a two-layer neural network and exhibits a separation of timescales and different learning phases whose respective sizes depends the timescale parameter $\gamma$. In [marion2023leveraging], the authors study two-timescale gradient descent for a simple model of $1$-dimensional neural network and show that the teacher network is recovered as soon as both the number of neurons of the student and the timescale parameter are sufficiently large. In [bietti2023learning], a multi-index regression problem is considered. Relying on the assumption of high dimensional gaussian data, the authors consider a linear layer composed with a non-parameteric model whose projection can be computed in the Hermite basis. They show this VarPro algorithm results in a saddle-to-saddle dynamic on the linear layer and establish guarantees for the recovery of the teacher model.
Finally, [takakura2024mean, wang2024mean] study the training of mean-field models of neural networks in the two-timescale limit with noisy gradient descent. In contrast with these works, we do not consider here additional entropic or $L^{2}$-regularization on the feature weights.
Wasserstein gradient flow of statistical distances
Under our Assumption 1, eq. 18 shows $\mathcal{L}^{\lambda}$ is an infimal convolution of statistical divergences between the feature distribution $\mu$ and the teacher $\bar{\nu}$, interpolating between the $\chi^{2}$-divergence $\chi^{2}(\bar{\nu}|\mu)$ (or more generally a $f$-divergence $\mathrm{D}_{f}(\bar{\nu}|\mu)$) when $\lambda\to 0^{+}$ and a (squared) kernel discrepancy $\mathrm{MMD}(\bar{\nu},\mu)^{2}$ when $\lambda\to\infty$. In the case $\lambda\to\infty$, gradient flows of $\mathrm{MMD}$-discrepancies and applications to sampling were studied in several works [arbel2019maximum, sejdinovic2013equivalence, hertrich2023wasserstein, hertrich2023generative, hertrich2024wasserstein, boufadene2023global]. Those flows are known to get trap in local minima but discrepancies associated to non-smooth kernel have been observed to behave better in terms of convergence [hertrich2023generative, hertrich2024wasserstein]. In the case of the coulomb kernel, [boufadene2023global] proves that the discrepancy loss admits no spurious local minima and that the discrepancy flow converges towards the target measure under regularity assumptions.
[…]
1.3 Mathematical preliminaries and notations
In the following, $\Omega$ will either be the $n$-dimensional Torus or a closed bounded convex domain of $\mathbb{R}^{n}$, for some $n\geq 1$. We note $\mathcal{M}(\Omega)$ the set of finite Borel measure over $\Omega$ and $\mathcal{P}(\Omega)$ is the subset of $\mathcal{M}(\Omega)$ consisting of probability measures. In particular, we will denote by $\pi\in\mathcal{P}(\Omega)$ the uniform distribution over $\Omega$. For a measure $\nu\in\mathcal{M}(\Omega)$, $|\nu|$ is its total-variation measure and $\|\nu\|_{\mathrm{TV}}$ is the total variation of $\nu$. For $p\in[1,+\infty)$, we note $\mathcal{W}_{p}$ the Wasserstein-$p$ distance defined for two probability measures $\mu,\mu^{\prime}\in\mathcal{P}(\Omega)$ by:
$\mathcal{W}_{p}(\mu,\mu^{\prime})\coloneqq\min_{\gamma\in\Gamma(\mu,\mu^{\prime})}\left(\int_{\Omega\times\Omega}\|\omega-\omega^{\prime}\|^{2}\mathrm{d}\gamma(\omega,\omega^{\prime})\right)^{1/p}$
[…]
For an integer $k\geq 0$ and $s\in(0,1]$, we note $\mathcal{C}^{k,s}(\Omega)$ (or just $\mathcal{C}^{k,s}$ when no ambiguity) the Hölder space of $k$-times continuously differentiable functions over $\Omega$ with $s$-Hölder $k^{\text{th}}$-derivative. We note .$\|.\|_{\mathcal{C}^{k,s}}$ the Hölder norm on $\mathcal{C}^{k,s}(\Omega)$. For a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$ and $p\in[1,+\infty]$, we note $L^{p}(\rho,\mathcal{C}^{k,s})$ the space of measurable functions $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$ s.t. $\phi(.,x)\in\mathcal{C}^{k,s}(\Omega)$ for $\mathrm{d}\rho$-a.e. $x\in\Omega$ and $\|\phi\|_{L^{p}(\rho,\mathcal{C}^{k,s})}\coloneqq\left(\int_{\mathbb{R}^{d}}\|\phi(.,x)\|^{p}_{\mathcal{C}^{k,s}}\mathrm{d}\rho(x)\right)^{1/p}<+\infty$. We will often use that if $\phi\in L^{2}(\rho,\mathcal{C}^{k,s})$ and $\alpha\in L^{2}(\rho)$ then the Böchner integral $\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x)$ is in $\mathcal{C}^{k,s}$ with:
$\left\|\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x)\right\|_{\mathcal{C}^{k,s}}\leq\|\phi\|_{L^{2}(\rho,\mathcal{C}^{k,s})}\|\alpha\|_{L^{2}(\rho)}.$
2 Reduced risk associated to the VarPro algorithm
We study in this work a VarPro algorithm or two-timescale limit of gradient descent for the training of neural networks. This strategy amounts to perform gradient descent on the reduced risk defined as the result of a partial minimization on a regularized version of the risk.
2.1 Primal formulation of the reduced risk
Whereas regularizing the risk with the euclidean square norm of the weights is a popular practice, the variable projection procedure can be used with other kind of regularization. Generally, for a convex function $f:\mathbb{R}\to\mathbb{R}$ and a regularization strength $\lambda>0$ we consider for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$:
$\mathcal{R}^{\lambda}_{f}(\mu,u)\coloneqq\frac{1}{2}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}+\lambda\int_{\Omega}f(u)\mathrm{d}\mu,$
where we assume $\mathcal{R}^{\lambda}_{f}(\mu,u)=+\infty$ if $f(u)$ is not integrable w.r.t. $\mu$. As before we consider the reduced risk obtained by minimizing $\mathcal{R}^{\lambda}_{f}$ w.r.t. the outer weights $u$. For every $\mu\in\mathcal{P}(\Omega)$ we define:
$\mathcal{L}^{\lambda}_{f}(\mu)\coloneqq\min_{u\in L^{1}(\mu)}\frac{1}{\lambda}\mathcal{R}^{\lambda}_{f}(\mu,u)=\min_{u\in L^{1}(\mu)}\frac{1}{2\lambda}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}+\int_{\Omega}f(u)\mathrm{d}\mu$ (9)
and this definition extends to the limiting case $\lambda\to 0^{+}$ by considering:
$\mathcal{L}^{0}_{f}(\mu)\coloneqq\min_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu.$ (10)
In the following, we always assume that $\phi\in L^{2}(\rho,\mathcal{C}^{0}(\Omega))$ (Assumption 3). This in particular implies that, for any $\mu\in\mathcal{P}(\Omega)$, the map $\Phi_{\mu}:L^{1}(\mu)\to L^{2}(\rho)$ is weakly continuous. We also consider the following assumptions on the regularization function:
Assumption 2.
The function $f:\mathbb{R}\to\mathbb{R}\cup{+\infty}$ is nonnegative, strictly convex and superlinear i.e. such that $\lim_{\pm\infty}\frac{f(t)}{|t|}=+\infty$.
By Lemma 2.1, this is sufficient to ensure the existence of a unique minimizer $u^{\lambda}_{f}[\mu]\in\operatorname*{arg\,min}\mathcal{R}^{\lambda}_{f}(\mu,u)$ when $\lambda>0$ and $u^{0}_{f}[\mu]\in\operatorname*{arg\,min}_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu$ when $\lambda=0$. Of particular interest in this work and more precisely in section 4 will be the case where $f(t)=|t|^{r}/(r-1)$ for some $r>1$. In this case we will note the corresponding reduced risk $\mathcal{L}^{\lambda}_{r}$. In particular, for $r=2$ we recover the “$L^{2}$-regularized” reduced risk defined in eq. 4 and eq. 5.
Lemma 2.1.
Assume Assumption 2 holds. Then, for every $\mu\in\mathcal{P}(\Omega)$, the functional
$\mathcal{I}_{f}:u\in L^{1}(\mu)\mapsto\int_{\Omega}f(u)\mathrm{d}\mu$
is strictly convex, weakly lower-semicontinuous and has weakly compact sublevel sets. In particular, eq. 9 (and eq. 10 if feasible) admits a unique minimizer $u^{\lambda}_{f}[\mu]$.
Proof.
Clearly $\mathcal{I}_{f}$ is strictly convex. Weak lower-semicontinuity is a classical consequence of the fact that $\mathcal{I}_{f}$ is convex and strongly lower-semicontinuous (using Fatou’s lemma), hence its epigraph is convex and strongly closed and hence also weakly closed. For weak compacity of sublevel sets, if $(u_{n})_{n\geq 0}$ is a sequence s.t. $\int_{\Omega}f(u_{n})\mathrm{d}\mu\leq C$ for every $n\geq 0$, then using that $f$ has super-linear growth, for every $\varepsilon>0$ there exists a $T\geq 0$ s.t. $|t|\leq\varepsilon f(t)$ for every $|t|\geq T$ and for every $n\geq 0$:
$\int_{|u_{n}|\geq T}|u_{n}|\mathrm{d}\mu\leq\varepsilon\int f(u_{n})\mathrm{d}\mu\leq\varepsilon C.$
Thus the sequence $(u_{n})_{n\geq 0}$ is uniformly integrable and admits a weakly converging subsequence by Dunford-Pettis theorem. ∎
2.2 Partial minimization on the space of measures
The reduced risk can also be obtain as the result of partial minimization of a convex functional over the space of measures. Whereas we have previously separated the role of the outer weights $u$ and of the feature distribution $\mu$ in eq. 2, our neural network model can equivalently be seen as a convolution of a measure on the feature space $\Omega$.
For $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$, we have by definition of $\Phi\star$ in eq. 7 and of $\Phi_{\mu}$ in eq. 3 that $\phi_{\mu}\cdot u=\Phi\star\nu$ where $\nu\in\mathcal{M}(\Omega)$ is s.t. $\mathrm{d}\nu=u\mathrm{d}\mu$. Also, $\int_{\Omega}f(u)\mathrm{d}\mu=\int_{\Omega}f(\frac{\mathrm{d}\nu}{\mathrm{d}\mu})\mathrm{d}\mu=\mathrm{D}_{f}(\nu|\mu)$ where, for $f$ satisfying Assumption 2, $\mathrm{D}_{f}$ is the divergence defined by:
$\forall(\nu,\mu)\in\mathcal{M}(\Omega)\times\mathcal{P}(\Omega),\quad\mathrm{D}_{f}(\nu|\mu)\coloneqq\left\{\begin{array}[]{cc}\int_{\Omega}f(\frac{\mathrm{d}\nu}{\mathrm{d}\mu})\mathrm{d}\mu&\text{if $\nu\ll\mu$},\\ +\infty&\text{otherwise}.\end{array}\right.$ (13)
In particular, in the case where $f$ is an entropy function and $\nu\in\mathcal{P}(\Omega)$ is a probability measure, $\mathrm{D}_{f}(\nu|\mu)$ is the standard Csiszàr $f$-divergence [liero2018optimal]. Performing a change of variable, one can thus define the functional $\mathcal{L}^{\lambda}_{f}$ as the value resulting from a minimization problem over the space of measures. For $\mu\in\mathcal{P}(\Omega)$, minimizing over $\nu\in\mathcal{M}(\Omega)$ instead of $u\in L^{1}(\mu)$, we get:
$\mathcal{L}_{f}^{\lambda}(\mu)=\left\{\begin{array}[]{cc}\min\limits_{\nu\in\mathcal{M}(\Omega)}\frac{1}{2\lambda}\|\Phi\star\nu-Y\|^{2}+\mathrm{D}_{f}(\nu|\mu)&\text{if $\lambda>0$,}\\[10.0pt] \min\limits_{\nu\in\mathcal{M}(\Omega)}\iota_{\Phi\star\nu=Y}+\mathrm{D}_{f}(\nu|\mu)&\text{if $\lambda=0$.}\end{array}\right.$ (16)
As presented in Assumption 1, of particular interest is the case where the signal $Y$ itself can be exactly represented by a neural network, that is $Y=\Phi\star\bar{\nu}$, for some $\bar{\nu}\in\mathcal{M}(\Omega)$. Then in the case $\lambda=0$, using the injectivity of $\Phi\star$, $\bar{\nu}$ is the only feasible solution in eq. 16 and we obtain:
$\mathcal{L}^{0}_{f}(\mu)=\int_{\Omega}f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu})\mathrm{d}\mu=\mathrm{D}_{f}(\bar{\nu}|\mu)\,.$ (17)
In the case $\lambda>0$, $\mathcal{L}^{\lambda}_{f}$ can be interpreted as the infimal convolution between a Maximum Mean Discrepancy (MMD) and the divergence $\mathrm{D}_{f}$. Indeed, naturally associated to the data distribution $\rho\in\mathcal{P}(\mathbb{R}^{d})$ and to the feature map $\phi$ is a structure of Reproducing Kernel Hilbert Space (RKHS) of functions that we will note $\mathcal{H}$ with a kernel $\kappa:\Omega\times\Omega\to\mathbb{R}$ defined by:
$\forall\omega,\omega^{\prime}\in\Omega,\quad\kappa(\omega,\omega^{\prime})\coloneqq\int_{\Omega}\phi(\omega,x)\phi(\omega^{\prime},x)\mathrm{d}\rho(x).$
We refer to appendix A for results of the theory of RKHSs we will use in this paper. Importantly, it follows from the definition of $\kappa$ and $\mathcal{H}$ that, under Assumption 1, the data attachment term in eq. 16 can be interpreted as a kernel distance between $\nu$ and $\bar{\nu}$. By eq. 43 we have $\|\Phi\star(\nu-\bar{\nu})\|_{L^{2}(\rho)}=\mathrm{MMD}_{\kappa}(\nu,\bar{\nu})$ where $\mathrm{MMD}_{\kappa}$ is the Maximum Mean Discrepancy (MMD) with kernel $\kappa$ [muandet2017kernel, gretton2012kernel]. For $\lambda>0$, the functional $\mathcal{L}^{\lambda}_{f}$ can then be expressed for every $\mu\in\mathcal{P}(\Omega)$ as:
$\mathcal{L}^{\lambda}_{f}(\mu)=\min_{\nu\in\mathcal{M}(\Omega)}\frac{1}{2\lambda}\mathrm{MMD}_{\kappa}^{2}(\nu,\bar{\nu})+\mathrm{D}_{f}(\nu|\mu).$ (18)
[…]
2.3 Dual formulation of the reduced risk
In eqs. 9 and 10, the objectives $\mathcal{L}^{\lambda}_{f}$ and $\mathcal{L}_{f}^{0}$ are expressed as a the value of a minimisation problem over the outer weights $u$. Taking the dual of those minimization problems, $\mathcal{L}_{f}^{\lambda}$ and $\mathcal{L}^{0}_{f}$ can be expressed as the value of a maximisation problem over the dual variable $\alpha\in L^{2}(\rho)$. In contrast with the primal formulation eq. 9, the dual formulation of Proposition 2.1 has the advantage of conveniently expressing $\mathcal{L}^{\lambda}_{f}$ for both $\lambda>0$ and $\lambda=0$ as the value of an optimization problem over the space $L^{2}(\rho)$ which is independent of $\mu$.
Proposition 2.1 (Dual representation).
Let Assumption 2 hold and consider $\mu\in\mathcal{P}(\Omega)$. Then we have for $\lambda>0$:
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)},$ (19)
where $f^{*}$ is the Legendre transform of $f$ and $\Phi^{\top}:L^{2}(\rho)\to\mathcal{C}^{0}(\Omega)$ is defined by:
$\forall\alpha\in L^{2}(\rho),\quad\Phi^{\top}\alpha=\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x).$
The supremum in eq. 19 is attained at some $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ and for $u^{\lambda}_{f}[\mu]\in L^{1}(\mu)$ the optimizer in eq. 9 it holds:
$\lambda\alpha^{\lambda}_{f}[\mu]=\Phi_{\mu}\cdot u^{\lambda}_{f}[\mu]-Y\quad\text{and}\quad f(u^{\lambda}_{f}[\mu])+f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu])=u^{\lambda}_{f}[\mu](\Phi^{\top}\alpha^{\lambda}_{f}[\mu]).$ (20)
Moreover, eq. 19 also holds in the case $\lambda=0$ under Assumption 1.
When $\lambda>0$, this result yields a convenient reformulation of the functional $\mathcal{L}^{\lambda}_{f}$. For $\mu\in\mathcal{P}(\Omega)$, $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ being the maximizer in eq. 19 and $u^{\lambda}_{f}[\mu]\in L^{1}(\mu)$ the minimizer in eq. 9, we have:
$\mathcal{L}^{\lambda}_{f}(\mu)=\frac{\lambda}{2}\left\|\alpha^{\lambda}_{f}[\mu]\right\|^{2}_{L^{2}(\rho)}+\int_{\Omega}f(u^{\lambda}_{f}[\mu])\mathrm{d}\mu.$ (21)
Proof.
Consider $\mu\in\mathcal{P}(\Omega)$ and $\lambda>0$. First, by definition of $\Phi^{\top}$ we have for every $\alpha\in L^{2}(\rho)$ and every $u\in L^{1}(\mu)$ that $\int_{\Omega}(\Phi^{\top}\alpha)u\mathrm{d}\mu=\langle\alpha,\Phi_{\mu}\cdot u\rangle_{L^{2}(\rho)}$ i.e. $\Phi^{\top}$ is the adjoint of $\Phi_{\mu}:L^{1}(\mu)\to L^{2}(\mu)$. Also, it follows from the assumption on $f$ that the map $\mathcal{I}_{f}:u\in L^{1}(\mu)\mapsto\int_{\Omega}f(u)\mathrm{d}\mu$ is a convex, weakly lower-semicontinuous functional whose Legendre transform is given for $h\in L^{\infty}(\mu)$ by:
$\mathcal{I}_{f}^{*}(h)=\sup_{u\in L^{1}(\mu)}\int_{\Omega}hu\mathrm{d}\mu-\int_{\Omega}f(u)\mathrm{d}\mu=\int_{\Omega}f^{*}(h)\mathrm{d}\mu,$
with $f^{*}$ the Legendre transform of $f$ and where the supremum is attained for $u\in L^{1}(\mu)$ satisfying the duality relation $f(u)+f^{*}(h)=uh$ [rockafellar1968integrals, Thm. 2]. Similarly, for $u\in L^{1}(\mu)$ we have:
$\sup_{\alpha\in L^{2}(\rho)}-\langle\alpha,\Phi_{\mu}\cdot u-Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)}=\frac{1}{2\lambda}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}.$
where the supremum is reached at $\alpha=\lambda\left(\Phi_{\mu}\cdot u-Y\right)$ when $\lambda>0$. Moreover, the functional $\alpha\mapsto\frac{1}{2\lambda}\|\alpha\|^{2}_{L^{2}(\rho)}$ being continuous, we can apply [rockafellar1967duality, Thm. 3] and eq. 19 holds by strong duality. The optimums are attained in both eq. 9 and eq. 19 and thus eq. 20 expresses the optimality conditions.
Finally, for the case $\lambda=0$, when Assumption 1 holds we have by eq. 16 that $\mathcal{L}^{0}_{f}(\mu)=\mathrm{D}_{f}(\bar{\nu}|\mu)$. Also, the assumptions on $f$ ensures $\mathrm{dom}(f^{*})=\mathbb{R}$ and using [rockafellar1971integrals, Thm. 4] we obtain:
$\mathcal{L}^{\lambda}_{f}(\mu)=\mathrm{D}_{f}(\bar{\nu}|\mu)=\sup_{h\in\mathcal{C}^{0}(\Omega)}\int_{\Omega}h\mathrm{d}\bar{\nu}-\int_{\Omega}f^{*}(h)\mathrm{d}\mu.$
The result follows as the injectivity of $\Phi\star$ ensures $\operatorname{Range}(\Phi^{\top})$ is dense in $\mathcal{C}^{0}(\Omega)$ (Lemma A.1).
Observing that $\Phi^{\top}$ defines a partial isometry from $L^{2}(\rho)$ to the RKHS $\mathcal{H}$ (eq. 41), a similar dual formulation of $\mathcal{L}^{\lambda}_{f}$ also holds in duality with $\mathcal{H}$.
Proposition 2.2.
Let Assumption 2 and Assumption 1 hold and consider $\mu\in\mathcal{P}(\Omega)$. Then we have for $\lambda\geq 0$:
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{h\in\mathcal{H}}-\int_{\Omega}f^{*}(h)\mathrm{d}\mu+\int_{\Omega}h\mathrm{d}\bar{\nu}-\frac{\lambda}{2}\|h\|^{2}_{\mathcal{H}},$ (22)
where $f^{*}$ is the Legendre transform of $f$. For $\lambda>0$, the supremum in eq. 22 is attained at some $h^{\lambda}_{f}[\mu]\in\mathcal{H}$ and for $\nu^{\lambda}_{f}[\mu]\in L^{1}(\mu)$ the optimizer in eq. 18 it holds:
$\lambda h^{\lambda}_{f}[\mu]=\Phi^{\top}\Phi\star(\nu^{\lambda}_{f}[\mu]-\bar{\nu})\quad\text{and}\quad f(\frac{\mathrm{d}\nu^{\lambda}_{f}[\mu]}{\mathrm{d}\mu})+f^{*}(h^{\lambda}_{f}[\mu])=h^{\lambda}_{f}[\mu]\frac{\mathrm{d}\nu^{\lambda}_{f}[\mu]}{\mathrm{d}\mu}.$ (23)
Proof.
The formula eq. 22 is directly deduced from eq. 19 and the characterization of the RKHS $\mathcal{H}$ in eq. 41. Also eq. 23 is a rewritting of eq. 20 since $\nu^{\lambda}_{f}[\mu]\in\mathcal{M}(\Omega)$ and $h^{\lambda}_{f}[\mu]\in\mathcal{H}$ are related to $u^{\lambda}_{f}[\mu]\in L^{1}(\mu)$ and $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ by $\mathrm{d}\nu^{\lambda}_{f}[\mu]=u^{\lambda}_{f}[\mu]\mathrm{d}\mu\quad\text{and}\quad h^{\lambda}_{f}[\mu]=\Phi^{\top}\alpha^{\lambda}_{f}[\mu]$. ∎
2.4 Kernel learning in the case of quadratic regularization
The case of a quadratic regularization is of particular interest since the partial optimization problem over $u$ admits a closed-form solution which can be efficiently obtained numerically by solving a linear system. In this case, the task of minimizing the reduced risk is equivalent to solving a Multiple Kernel Learning problem [bach2004multiple].
For the $L^{2}$-regularization $f(t)=|t|^{2}$, the reduced risk $\mathcal{L}^{\lambda}_{2}[\mu]$ is the value of the ridge regression problem in eq. 4 and for $\lambda>0$, the optimizer is given by $u^{\lambda}_{2}[\mu]=(\Phi_{\mu}^{\top}\Phi_{\mu}+2\lambda)^{-1}\Phi_{\mu}^{\top}Y$ where $\Phi_{\mu}^{\top}:L^{2}(\rho)\to L^{2}(\mu)$ is the adjoint of the operator $\Phi_{\mu}$ restricted to $L^{2}(\mu)$. Also, the dual problem reads:
$\mathcal{L}^{\lambda}_{2}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\frac{1}{2}\langle\alpha,(K_{\mu}+2\lambda)\alpha\rangle_{L^{2}(\rho)}+\langle\alpha,Y\rangle_{L^{2}(\rho)},$
where $K_{\mu}:L^{2}(\rho)\to L^{2}(\rho)$ is the self-adjoint operator defined by $K_{\mu}=\Phi_{\mu}\Phi_{\mu}^{\top}$. The supremum is attained at $\alpha^{\lambda}_{2}[\mu]=(K_{\mu}+2\lambda)^{-1}Y$ and by eq. 21 we obtain for every $\mu\in\mathcal{P}(\Omega)$:
$\mathcal{L}^{\lambda}_{2}(\mu)=\frac{1}{2}\langle Y,(K_{\mu}+2\lambda)^{-1}Y\rangle_{L^{2}(\rho)}.$ (24)
This is the optimal value of the kernel ridge regression problem with kernel $K_{\mu}$ parameterized by the feature distribution $\mu$. Moreover this parameterization is linear w.r.t. $\mu\in\mathcal{P}(\Omega)$ since considering for $\omega\in\Omega$ the rank the rank-one self-adjoint operator $k(\omega)\coloneqq\phi(\omega,.)\otimes\phi(\omega,.)$ we have:
$K_{\mu}=\int_{\Omega}k(\omega)\mathrm{d}\mu(\omega).$
Therefore, minimizing the reduced risk $\mathcal{L}^{\lambda}_{2}$ over the feature distribution $\mu$ amounts to find the best kernel for solving the ridge regression problem in eq. 4 among convex combinations of “simple” basis kernels $(k(\omega))_{\omega\in\Omega}$ i.e. a Multiple Kernel Learning task. Other convex optimization strategies for solving such task have been studied in [lanckriet2004learning, bach2004multiple].
3 Properties of minimizers of the reduced risk
Before turning to the analysis of gradient methods for the minimization of the reduced risk $\mathcal{L}^{\lambda}_{f}$ in sections 4 and 5, we study here variational properties of $\mathcal{L}^{\lambda}_{f}$.
3.1 Existence and uniqueness of minimizers
We first investigate existence and uniqueness of minimizers of $\mathcal{L}^{\lambda}_{f}$. Importantly, we use here that $\mathcal{L}^{\lambda}_{f}$ is obtained as a the result of a partial minimization. Namely, for $\lambda\geq 0$ and $\mu\in\mathcal{P}(\Omega)$, we have from eq. 16 that $\mathcal{L}^{\lambda}_{f}(\mu)=\min_{\nu\in\mathcal{M}(\Omega)}\mathcal{E}^{\lambda}_{f}(\nu,\mu)$, where $\mathcal{E}^{\lambda}_{f}$ is defined for $\nu\in\mathcal{M}(\Omega)$ and $\mu\in\mathcal{P}(\Omega)$ by:
$\mathcal{E}^{\lambda}_{f}(\nu,\mu)\coloneqq\left\{\begin{array}[]{cc}\mathrm{D}_{f}(\nu|\mu)+\frac{1}{2\lambda}\|\Phi\star\nu-Y\|^{2}&\text{if $\lambda>0$,}\\ \mathrm{D}_{f}(\nu|\mu)+\iota_{\Phi\star\nu=Y}&\text{if $\lambda=0$.}\end{array}\right.$ (27)
In particular, it follows from variational formulations of the $f$-divergences that $\mathrm{D}_{f}$ is (jointly) convex and lower-semicontinuous w.r.t. its arguments $(\nu,\mu)\in\mathcal{M}(\Omega)\times\mathcal{P}(\Omega)$ [rockafellar1971integrals, Thm. 4]. The following Lemma 3.1 uses this fact to establish convexity and lower-semicontinuity of $\mathcal{L}^{\lambda}_{f}$, implying the existence of minimizers. We then discuss cases in which $\mathcal{L}^{\lambda}_{f}$ has in fact a unique minimizer.
Lemma 3.1.
Assume $f$ satisfies Assumption 2. Then, for $\lambda\geq 0$, $\mathcal{L}^{\lambda}_{f}:\mathcal{P}(\Omega)\to\mathbb{R}$ is a convex, lower semi-continuous function (w.r.t. the narrow convergence on $\mathcal{M}(\Omega)$).
Proof.
By the definition of the divergence $\mathrm{D}_{f}$ in eq. 13 and by [rockafellar1971integrals, Thm. 4], we have for every $(\nu,\mu)\in\mathcal{M}(\Omega)\times\mathcal{P}(\Omega)$:
$\mathrm{D}_{f}(\nu|\mu)=\sup_{h\in\mathcal{C}^{0}(\Omega)}\int_{\Omega}h\mathrm{d}\nu-\int_{\Omega}f^{*}(h)\mathrm{d}\mu.$
Thus $\mathrm{D}_{f}$ is a (jointly) convex and lower-semicontinuous function as a supremum of (jointly) convex and lower semi-continuous functions. As a consequence, for $\lambda\geq 0$, $\mathcal{E}^{\lambda}_{f}$ is also (jointly) convex and lower-semicontinuous and the convexity of $\mathcal{L}^{\lambda}_{f}=\min_{\nu}\mathcal{E}^{\lambda}_{f}(\nu,.)$ follows as partial minimization preserves convexity. Also if $(\mu_{n})_{n\geq 0}$ is a sequence in $\mathcal{P}(\Omega)$ converging narrowly to some $\mu\in\mathcal{P}(\Omega)$, then we have $\mathcal{L}^{\lambda}_{f}(\mu_{n})=\mathcal{E}^{\lambda}_{f}(\nu_{n},\mu_{n})$ for some $\nu_{n}\in\mathcal{M}(\Omega)$. Without loss of generality one can assume $\mathcal{L}^{\lambda}_{f}(\mu_{n})$ is bounded, thus $\mathrm{D}_{f}(\nu_{n}|\mu_{n})$ and then $\|\nu_{n}\|_{\mathrm{TV}}$ are bounded as $f$ is superlinear. Then, up to extraction of a subsequence, $(\nu_{n})$ converges narrowly to $\nu\in\mathcal{M}(\Omega)$ and we get by lower semi-continuity of $\mathcal{E}^{\lambda}_{f}$:
$\liminf_{n\to\infty}\mathcal{L}^{\lambda}_{f}(\mu_{n})=\liminf_{n\to\infty}\mathcal{E}^{\lambda}_{f}(\nu_{n},\mu_{n})\geq\mathcal{E}^{\lambda}_{f}(\nu,\mu)\geq\mathcal{L}^{\lambda}_{f}(\mu),$
which shows that $\mathcal{L}^{\lambda}_{f}$ is lower-semicontinuous.
The above result implies the existence of minimizers of the reduced risk $\mathcal{L}^{\lambda}_{f}$ for every $\lambda\geq 0$ but it does not establish uniqueness and $\mathcal{L}^{\lambda}_{f}$ may, a priori, have several minimizers. However, there are cases in which uniqueness can be ensured, we give two examples:
  • In the teacher student setup where Assumption 1 holds, if $\bar{\nu}$ is a positive measure with $\bar{m}=\bar{\nu}(\Omega)>0$ and if $f$ is nonnegative, strictly convex and such that $f(\bar{m})=0$ then the teacher feature distribution $\bar{\mu}=\bar{\nu}/\bar{m}$ is the unique minimizer of $\mathcal{L}^{\lambda}_{f}$, whatever $\lambda\geq 0$. Indeed, from eq. 18 we have that $\mathcal{L}^{\lambda}_{f}(\bar{\mu})=0$ and $\mathcal{L}^{\lambda}_{f}(\mu)>0$ for every $\mu\neq\bar{\mu}$. With these assumptions, we prove in Theorem 5.1 that the gradient flow of $\mathcal{L}^{\lambda}_{f}$ converges towards the teacher feature distribution $\bar{\mu}$ with an algebraic convergence rate.
  • For general data $Y$, relying on a variational characterization of the total variation, the following Lemma 3.2 establishes uniqueness of a minimizer to $\mathcal{L}^{\lambda}_{r}$ in the case the regularization is of the form $f(t)=|t|^{r}$ for some $r>1$.
Lemma 3.2.
Let $\lambda\geq 0$ and assume $f(t)=|t|^{r}$ for some $r>1$. Then $\mathcal{L}^{\lambda}_{r}$ admits a unique minimizer $\bar{\mu}^{\lambda}_{r}\in\mathcal{P}(\Omega)$.
Proof.
We use arguments similar to the one of [wang2024mean, Prop. 3.3]. By duality $\inf_{\mu\in\mathcal{P}(\Omega)}\mathcal{L}^{\lambda}_{r}(\mu)=\inf_{\nu\in\mathcal{M}(\Omega)}\mathcal{G}^{\lambda}_{r}(\nu)$ where $\mathcal{G}^{\lambda}_{r}$ is defined for $\forall\nu\in\mathcal{M}(\Omega)$ by:
$\mathcal{G}^{\lambda}(\nu)\coloneqq\left\{\begin{array}[]{cc}\|\nu\|^{r}_{\mathrm{TV}}+\frac{1}{2\lambda}\|\Phi\star\nu-Y\|^{2}&\text{if $\lambda>0$,}\\ \|\nu\|^{r}_{\mathrm{TV}}+\iota_{\Phi\star\nu=Y}&\text{if $\lambda=0$,}\end{array}\right.$
where we used the variational representation $\|\nu\|^{r}_{\mathrm{TV}}=\inf_{\mu\in\mathcal{P}(\Omega)}\int_{\Omega}\left|\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu$ (the case $r=2$ is used in [wang2024mean, lanckriet2004learning]). Indeed, for measures $\nu\in\mathcal{M}(\Omega)$ and $\mu\in\mathcal{P}(\Omega)$ s.t. $\nu\ll\mu$ we have:
$\int_{\Omega}\left|\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu=\int_{\Omega}\left(\frac{\mathrm{d}\mu}{\mathrm{d}|\nu|}\right)^{1-r}\mathrm{d}|\nu|.$
The Lagrangian of the convex problem $\inf_{\mu}\int_{\Omega}\left(\frac{\mathrm{d}\mu}{\mathrm{d}|\nu|}\right)^{1-r}\mathrm{d}|\nu|$ is given by:
$\mathcal{J}(\mu,\gamma)=\int_{\Omega}\left(\frac{\mathrm{d}\mu}{\mathrm{d}|\nu|}\right)^{1-r}\mathrm{d}|\nu|+\gamma\left(\int_{\Omega}\mathrm{d}\mu-1\right).$
The optimality condition gives that $\frac{\mathrm{d}\mu}{\mathrm{d}|\nu|}$ is constant and the minimum is attained for $\mu=|\nu|/\|\nu\|_{\mathrm{TV}}$, giving $\int_{\Omega}\left|\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu=\|\nu\|^{r}_{\mathrm{TV}}$. Finally, the map $\nu\mapsto\|\nu\|^{r}_{\mathrm{TV}}$ is strictly convex so that $\mathcal{G}^{\lambda}_{r}$ admits a unique minimizer $\bar{\nu}^{\lambda}_{r}\in\mathcal{M}(\Omega)$, thus $\mathcal{L}^{\lambda}_{r}$ has also a unique minimizer $\bar{\mu}^{\lambda}_{r}\in\mathcal{P}(\Omega)$ and we have the duality relation $\bar{\mu}^{\lambda}_{r}=|\bar{\nu}^{\lambda}_{r}|/\|\bar{\nu}^{\lambda}_{r}\|_{\mathrm{TV}}$. Notably, in the case where Assumption 1 holds and $\lambda=0$, for every $r>1$, we have $\bar{\nu}^{0}_{r}=\bar{\nu}$ and $\bar{\mu}^{0}_{r}=\bar{\mu}$. ∎
3.2 Convergence of minimizers
Of particular interest to us is the case $\lambda=0$ for which minimizers of $\mathcal{L}^{0}_{f}$ are related to the teacher measure $\bar{\nu}$ by eq. 17. However in practice, minimization of $\mathcal{L}^{\lambda}_{f}$ is easier in the presence of a regularization parameter $\lambda>0$. For this reason, we are interested in the asymptotic behavior of minimizers of $\mathcal{L}^{\lambda}_{f}$ when $\lambda\to 0^{+}$.
[…]
Lemma 3.3 ($\Gamma$-convergence).
Assume Assumptions 2 and 1 hold, $\phi\in L^{2}(\rho,\mathcal{C}^{0,1})$ and $f^{*}\in\mathcal{C}^{0,1}_{\mathrm{loc}}(\mathbb{R})$. Then the family of functionals $(\mathcal{L}^{\lambda}_{f})_{\lambda>0}$ $\Gamma$-converges towards $\mathcal{L}^{0}_{f}$ as $\lambda\to 0^{+}$ in the sense that for every sequence $(\lambda_{n})_{n\geq 0}$ converging to $0^{+}$ and every $\mu\in\mathcal{P}(\Omega)$ it holds:
  1. 1.
    for every sequence $(\mu_{n})_{n\geq 0}$ converging narrowly to $\mu$, $\liminf_{n\to\infty}\mathcal{L}^{\lambda_{n}}_{f}(\mu_{n})\geq\mathcal{L}^{0}_{f}(\mu)_{f}$,
  2. 2.
    there exists a sequence $(\mu_{n})_{n\geq 0}$ converging narrowly to $\mu$ s.t. $\limsup_{n\to\infty}\mathcal{L}^{\lambda_{n}}_{f}(\mu_{n})\leq\mathcal{L}^{0}_{f}(\mu)$.
Proof.
For the second part of the result it suffices to consider the constant sequence $\mu_{n}=\mu$ for every $n\geq 0$. Indeed, it then directly follows from the definition of $\mathcal{L}^{\lambda}_{f}$ and $\mathcal{L}^{0}_{f}$ in eq. 9 and eq. 10 that $\mathcal{L}^{\lambda}_{f}(\mu)\leq\mathcal{L}^{0}_{f}(\mu)$ for every $\lambda>0$.
To prove the first part of the definition, consider a sequence $(\mu_{n})_{n\geq 0}$ converging narrowly to $\mu$ in $\mathcal{P}(\Omega)$. By the dual formulation of $\mathcal{L}^{0}_{f}$ in eq. 19, we can also consider a sequence $(\alpha_{k})_{k\geq 0}$ in $L^{2}(\rho)$ such that:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\langle\alpha_{k},Y\rangle\xrightarrow{k\to\infty}\mathcal{L}^{0}_{f}(\mu).$
Then by the dual formulation of $\mathcal{L}^{\lambda}_{f}$ in eq. 19, for every $n,k\geq 0$:
$\mathcal{L}^{\lambda_{n}}(\mu_{n})$ $\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu_{n}+\langle\alpha_{k},Y\rangle-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}$
$=-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}(\mu_{n}-\mu)-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\langle\alpha_{k},Y\rangle\rangle$
$\geq-\left\|f^{*}(\Phi^{\top}\alpha_{k})\right\|_{\mathcal{C}^{0,1}}\mathcal{W}_{1}(\mu_{n},\mu)-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\langle\alpha_{k},Y\rangle\rangle.$
But then, since $\mathcal{W}_{1}(\mu_{n},\mu)\to 0$ and $\lambda_{n}\to 0$, one can find a sequence $(k_{n})_{n\geq 0}$ diverging to infinity such that it holds:
$\lambda_{n}\|\alpha_{k_{n}}\|^{2}_{L^{2}(\rho)}\xrightarrow{n\to+\infty}0\quad\text{and}\quad\left\|f^{*}(\Phi^{\top}\alpha_{k_{n}})\right\|_{\mathcal{C}^{0,1}}\mathcal{W}_{1}(\mu_{n},\mu)\xrightarrow{n\to+\infty}0.$
Thus, for every $n\geq 0$, $\mathcal{L}^{\lambda_{n}}_{f}(\mu_{n})\geq\mathcal{L}^{0}_{f}(\mu)+o(1)$ and the result follows. ∎
It is a direct consequence of the above $\Gamma$-convergence result that the limit when $\lambda\to 0^{+}$ of a sequence of minimizers of $\mathcal{L}^{\lambda}_{f}$ is a minimizer of $\mathcal{L}^{0}_{f}$ [santambrogio2023course, Prop. 7.5]. In the case where $\mathcal{L}^{\lambda}_{f}$ has a unique minimizer $\bar{\mu}^{0}_{f}$, this implies every sequence of minimizers of $\mathcal{L}^{\lambda}_{f}$ converges to $\bar{\mu}^{0}_{f}$ when $\lambda\to 0^{+}$.
Proposition 3.1 (Convergence of minimizers).
Assume the result of Lemma 3.3 holds. For every $\lambda>0$, consider $\bar{\mu}^{\lambda}_{f}\in\operatorname*{arg\,min}\mathcal{L}^{\lambda}_{f}$ and assume $\bar{\mu}^{\lambda}_{f}\xrightarrow{\lambda\to 0^{+}}\mu$. Then $\mu\in\operatorname*{arg\,min}\mathcal{L}^{0}_{f}$. Notably, if $\mathcal{L}^{0}_{f}$ has a unique minimizer $\bar{\mu}^{0}_{f}$, then $\bar{\mu}^{\lambda}_{f}\xrightarrow{\lambda\to 0^{+}}\bar{\mu}^{0}_{f}$.
4 Training with gradient flow
In the rest of this work we consider the optimization over the feature distribution $\mu\in\mathcal{P}(\Omega)$ for the minimization of the reduced risk $\mathcal{L}^{\lambda}_{f}$, for $\lambda\geq 0$. Specifically, we consider a gradient flow algorithm. In the case of a finite number of feature $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ such gradient flow is defined as the solution of the equation:
$\forall i\in\{1,...,M\},\quad\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)=-M\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}_{f}(\{\omega_{i}(t)\}_{1\leq i\leq M})$ (28)
where ^$\hat{\mathcal{L}}^{\lambda}_{f}(\{\omega_{i}\}_{1\leq i\leq M})\coloneqq\mathcal{L}^{\lambda}_{f}(\hat{\mu})$ and $\hat{\mu}$ is the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. More generally, in terms of the feature distribution $\mu\in\mathcal{P}(\Omega)$, the above equation corresponds to a Wasserstein gradient flow over the functional $\mathcal{L}^{\lambda}_{f}$, namely:
$\partial_{t}\mu_{t}-\mathrm{div}\left(\mu_{t}\nabla\frac{\delta\mathcal{L}^{\lambda}_{f}}{\delta\mu}[\mu_{t}]\right),\quad\text{over $(0,\infty)\times\Omega$,}$ (29)
[…]
$\forall k\geq 0,\quad\mu^{\tau}_{k+1}\in\operatorname*{arg\,min}_{\mu\in\mathcal{P}(\Omega)}\mathcal{L}^{\lambda}_{f}(\mu)+\frac{1}{2\tau}\mathcal{W}_{2}(\mu,\mu^{\tau}_{k})^{2}.$ (30)
We study in this section the well-posedness of the above Wasserstein gradient flow equation by distinguishing the case where $\lambda>0$ and the case $\lambda=0$. In the latter case, we show the Wasserstein gradient flow corresponds to an ultra-fast diffusion equation.
4.1 Wasserstein gradient flows in the case $\lambda>0$
In the case where $\lambda>0$, the presence of the regularization induces sufficient regularity on the objective to study the training dynamic through the lens of classical results from the theory of gradient flows in the Wasserstein space [ambrosio2008gradient, santambrogio2017euclidean]. In particular, one can derive the gradient flow equation leveraging the dual representation of $\mathcal{L}^{\lambda}_{f}$. Indeed, eq. 19 expresses $\mathcal{L}^{\lambda}_{f}$ as a maximum over linear functionals, and thus by the envelope theorem one can formally differentiate $\mathcal{L}^{\lambda}_{f}$ w.r.t. $\mu$ and obtain the Fréchet differential:
$\frac{\delta\mathcal{L}^{\lambda}_{f}}{\delta\mu}[\mu](\omega)=-f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu])(\omega)$
with $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ the maximizer in eq. 19. We show that the gradient field of this potential indeed defines a notion of “gradient” for the functional $\mathcal{L}^{\lambda}_{f}$ w.r.t. the Wasserstein-$2$ topology on $\mathcal{P}(\Omega)$. More precisely, absolutely continuous curves $(\mu_{t})_{t\in[0,1]}$ in the space $\mathcal{P}(\Omega)$ are characterised as solutions to a continuity equation:
$\partial_{t}\mu_{t}+\mathrm{div}(\mu_{t}v_{t})=0\quad\text{over $(0,1)\times\Omega$}$ (31)
for some velocity field $v$ such that $\|v_{t}\|_{L^{2}(\mu_{t})}\in L^{1}((0,1))$ [santambrogio2015optimal, Thm. 5.14]. This equation has to be understood in the sense of distributions, that is in duality with the set $\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega)$ of compactly supported functions, i.e.:
$\int_{0}^{1}\int_{\Omega}\left(\partial_{t}\varphi+\langle\nabla\varphi,v_{t}\rangle\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,1)\times\Omega).$ (32)
The following result shows that the functional $\mathcal{L}^{\lambda}_{f}(\mu_{t})$ is differentiable along those curves and expresses its derivative in terms of the gradient field $\nabla\frac{\delta\mathcal{L}^{\lambda}_{f}}{\delta\mu}$.
Lemma 4.1 (Wasserstein chain rule for $\mathcal{L}^{\lambda}_{f}$).
Assume $\phi\in L^{2}(\rho,\mathcal{C}^{1})$, $f$ satisfies Assumption 2 with $f^{*}\in\mathcal{C}^{1}_{\mathrm{loc}}(\mathbb{R})$ and consider $\lambda>0$. Let $(\mu_{t})_{t\in[0,1]}$ be an absolutely continuous curve in $\mathcal{P}(\Omega)$ solution of the continuity equation eq. 32 for some velocity field $v$ such that $\|v_{t}\|_{L^{2}(\mu_{t})}\in L^{1}((0,1))$. Then $(\mathcal{L}^{\lambda}_{f}(\mu_{t}))_{t\in[0,1]}$ is absolutely continuous and for a.e. $t^{\prime},t\in[0,1]$:
$\mathcal{L}^{\lambda}_{f}(\mu_{t^{\prime}})-\mathcal{L}^{\lambda}_{f}(\mu_{t})=\int_{t}^{t^{\prime}}\langle\nabla\mathcal{L}^{\lambda}_{f}[\mu_{s}],v_{s}\rangle_{L^{2}(\mu_{s})}\mathrm{d}s$
where, for $\mu\in\mathcal{P}_{2}(\Omega)$, the velocity field $\nabla\mathcal{L}^{\lambda}_{f}[\mu]\in L^{2}(\mu)$ is defined as:
$\forall\omega\in\Omega,\quad\nabla\mathcal{L}^{\lambda}_{f}[\mu](\omega)\coloneqq-\nabla\left(f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu])\right)(\omega)$
with $\alpha^{\lambda}_{f}[\mu]$ the maximizer in eq. 19.
Proof.
Consider the dual formulation of $\mathcal{L}^{\lambda}_{f}$ in eq. 19. For every $\mu\in\mathcal{P}(\Omega)$ we have:
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}\mathcal{V}_{\alpha}(\mu)-\frac{\lambda}{2}\|\alpha\|^{2}+\langle\alpha,Y\rangle,$
where for $\alpha\in L^{2}(\rho)$ we defined:
$\mathcal{V}_{\alpha}(\mu)\coloneqq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)(\omega)\mathrm{d}\mu(\omega).$
In particular, at fixed $\alpha\in L^{2}(\rho)$, it follows from the assumptions on $\phi$ and $f^{*}$ that the potential $f^{*}(\Phi^{\top}\alpha)$ is in $\mathcal{C}^{1}(\Omega)$ with $\|f^{*}(\Phi^{\top}\alpha)\|_{\mathcal{C}^{1}}\leq C(\|\alpha\|_{L^{2}(\rho)})$ for some continuous function $C$. Thus, by properties of the continuity equation, $\mathcal{V}_{\alpha}(\mu_{t})$ is absolutely continuous and its derivative is give for a.e. $t\in[0,1]$ by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{V}_{\alpha}(\mu_{t})=-\int_{\Omega}\langle\nabla f^{*}(\Phi^{\top}\alpha),v_{t}\rangle\mathrm{d}\mu_{t}.$
Moreover, for $\mu\in\mathcal{P}(\Omega)$, using eq. 21 and the fact that $\mathcal{L}^{\lambda}_{f}\leq\frac{1}{2\lambda}\|Y\|^{2}_{L^{2}(\rho)}+f(0)$ (by taking $u=0$ in eq. 9) we have at the optimum in eq. 19 that
$\|\alpha^{\lambda}_{f}[\mu]\|_{L^{2}(\rho)}\leq\lambda^{-1}\left(\|Y\|^{2}_{L^{2}(\rho)}+\lambda f(0)\right)^{1/2}\eqqcolon R_{\lambda}$
Thus, $\mathcal{L}^{\lambda}_{f}$ is equivalently defined by restricting the supremum to $\alpha\in L^{2}(\rho)$ such that $\|\alpha\|_{L^{2}(\rho)}\leq R_{\lambda}$. For such $\alpha$ we have $\left|\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{V}_{\alpha}(\mu_{t})\right|\leq C^{\prime}$ for some constant $C^{\prime}=C^{\prime}(f^{*},\lambda)$ independent of $\alpha$. Thus we can apply the envelope theorem in [milgrom2002envelope, Thm. 2], which shows the desired result. ∎
The preceding result has defined a notion of gradient field for the functional $\mathcal{L}^{\lambda}_{f}$. One can thus define gradient flows of $\mathcal{L}^{\lambda}_{f}$ for the $\mathcal{W}_{2}$ metric as the curves solution to the continuity equation:
$\partial_{t}\mu_{t}-\mathrm{div}(\mu_{t}\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}])=0\quad\text{over $(0,\infty)\times\Omega$}.$ (33)
We make the following definition:
Definition 4.1 (Gradient flow of $\mathcal{L}^{\lambda}_{f}$).
Let $\mu_{0}\in\mathcal{P}_{2}(\Omega)$. We say $(\mu_{t})_{t\geq 0}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ starting at $\mu_{0}$ if it is a locally absolutely continuous curve s.t. $\lim_{t\to 0^{+}}\mu_{t}=\mu_{0}$ and satisfying the continuity equation eq. 33 in the sense of distribution, i.e.:
$\int_{0}^{\infty}\int_{\Omega}\left(\partial_{t}\varphi-\nabla\varphi\cdot\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega).$ (34)
Remark 4.1 (Boundary conditions).
Note that, in the case where $\Omega$ is a closed, bounded, smooth and convex domain, our definition eq. 32 of solutions to the continuity equation enforces no-flux conditions on the boundary $\partial\Omega$. Indeed we consider test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,1)\times\Omega)$ that can be supported on the whole domain $\Omega$ (which is always assumed closed). Thus, eq. 32 enforces $\langle\mu_{t}v_{t},\vec{n}\rangle=0$ in the sense of distribution, where $\vec{n}$ is the outer normal vector to the boundary $\partial\Omega$.
In case of the gradient flow equation eq. 34, this boundary condition is for example satisfied if one assumes $\langle\nabla_{\omega}\phi(\omega,x),\vec{n}\rangle=0$ for every $x\in\mathbb{R}^{d}$ and every $\omega\in\partial\Omega$. Another way of ensuring the no-flux condition is to remove the outer part of the gradient field $\nabla\mathcal{L}^{\lambda}_{f}$ on the boundary $\partial\Omega$, which can be performed by clipping the features.
Well-posedness of the gradient flow equation
To show the well-posedness of gradient flows, we rely on convexity properties of the functional $\mathcal{L}^{\lambda}_{f}$. Indeed, by the dual formulation in eq. 19, we can express $\mathcal{L}^{\lambda}_{f}$ as a supremum over semi-convex functional. As a consequence, the Lemma 4.2 below shows that, for $\lambda>0$, $\mathcal{L}^{\lambda}_{f}$ is semi-convex along (generalized) geodesics of the Wasserstein space (see [ambrosio2008gradient, Def. 9.2.4] for the definiton of generalized geodesics). However, note that such an argument can not be extended to the case $\lambda=0$ since the semi-convexity constant blows-up when $\lambda\to 0^{+}$. For example, in the case $f(t)=|t|^{2}$ this constant scales as $\lambda^{-2}$.
Lemma 4.2 (Geodesic semi-convexity).
Assume $\phi\in L^{2}(\rho,\mathcal{C}^{1,1})$, $f$ satisfies Assumption 2 with $f^{*}\in\mathcal{C}^{1,1}_{\mathrm{loc}}(\mathbb{R})$ and let $\lambda>0$. Then $\mathcal{L}^{\lambda}_{f}$ is $C$-semiconvex along (generalized) geodesics for some constant $C=C(f^{*},\lambda)$.
Proof.
Consider the dual formulation of $\mathcal{L}^{\lambda}_{f}$ in eq. 19. For every $\mu\in\mathcal{P}(\Omega)$ we have:
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)(\omega)\mathrm{d}\mu(\omega)-\frac{\lambda}{2}\|\alpha\|^{2}+\langle\alpha,Y\rangle,$
Then, at fixed $\alpha\in L^{2}(\rho)$, it follows from the assumptions on $\phi$ that $\Phi^{\top}\alpha\in\mathcal{C}^{1,1}(\Omega)$ with
$\|\Phi^{\top}\alpha\|_{\mathcal{C}^{1,1}}\leq\|\alpha\|_{L^{2}(\rho)}\|\phi\|_{L^{2}(\rho,\mathcal{C}^{1,1})}.$
Then, from the assumptions on $f^{*}$, the composition $f^{*}(\Phi^{\top}\alpha)$ is also in $\mathcal{C}^{1,1}(\Omega)$ and by [ambrosio2008gradient, Prop.9.3.2] the functional $\mu\mapsto\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu$ is $C$-semiconvex along generalized geodesic for some constant $C=C(f^{*},\|\alpha\|_{L^{2}(\rho)}\|\phi\|_{L^{2}(\rho,\mathcal{C}^{1,1})})$. Moreover, similarly as in the proof of Lemma 4.1, one can restrict the definition of $\mathcal{L}^{\lambda}_{f}$ to the supremum over $\alpha\in L^{2}(\rho)$ with $\|\alpha\|_{L^{2}(\rho)}\leq R_{\lambda}$. The result then follows by taking a supremum over (uniformly) semi-convex functionals. ∎
The semi-convexity of $\mathcal{L}^{\lambda}_{f}$ along generalized geodesic ensures the existence and uniqueness of gradient flows in the sense of Definition 4.1.
Theorem 4.1 (Well-posedness of the gradient flow equation for $\lambda>0$).
Assume the assumptions of Lemma 3.1 hold. Then for any $\lambda>0$ and any initialization $\mu_{0}\in\mathcal{P}_{2}(\Omega)$ there exists a unique gradient flow for $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ in the sense of Definition 4.1. Moreover, if $(\mu_{t})_{t\geq 0},(\mu_{t}^{\prime})_{t\geq 0}$ are gradient flows for $\mathcal{L}^{\lambda}_{f}$ with respective initializations $\mu_{0},\mu_{0}^{\prime}\in\mathcal{P}(\Omega)$ then for every $t\geq 0$:
$\mathcal{W}_{2}(\mu_{t},\mu_{t}^{\prime})\leq e^{Ct}\mathcal{W}_{2}(\mu_{0},\mu_{0}^{\prime}),$
for some constant $C=C(f^{*},\lambda)$.
Proof.
The chain rule formula established in Lemma 4.1 shows that for every $\mu\in\mathcal{P}(\Omega)$ the vector field $\mathcal{L}^{\lambda}_{f}[\mu]$ is a strong subdifferential of $\mathcal{L}^{\lambda}_{f}$ in the sense of [ambrosio2008gradient, Def. 10.3.1 and eq. (10.3.12)]. Existence, uniqueness and properties of the gradient flows then follows from the geodesic semi-convexity of $\mathcal{L}^{\lambda}_{f}$ established in Lemma 4.2 and the application of [ambrosio2008gradient, Def. 11.2.1]
Finally, it is a classical property of weak solutions to continuity equations that the gradient flows of $\mathcal{L}^{\lambda}_{f}$ can be represented in terms of push-forward by a flow map.
Proposition 4.1.
Assume the assumptions of Lemma 3.1 hold. Then, for any $\lambda>0$ and any initialization $\mu_{0}\in\mathcal{P}_{2}(\Omega)$, the gradient flow $(\mu_{t})_{t\geq 0}$ of $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ satisfies $\mu_{t}=(X_{t})_{\#}\mu_{0}$ for every $t\geq 0$, where $(X_{t})_{t\geq 0}$ is the flow-map solution of the ODE:
$\forall t\geq 0,\quad\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]\circ X_{t},\quad\text{with}\quad X_{0}=\mathrm{Id}_{\Omega}.$
In particular, if $(\omega_{i}(t))_{t\geq 0}$ for $i\in\{1,...,M\}$ are solutions to eq. 28 then the empirical distribution $\hat{\mu}_{t}\coloneqq\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}(t)}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ in the sense of Definition 4.1 and thus $\omega_{i}(t)=X_{t}(\omega_{i}(0))$ for $i\in\{1,...,M\}$ and $t\geq 0$.
Proof.
For every $t\geq 0$, similarly as in the proof of Lemma 4.1 we have that the dual variable is bounded by $\|\alpha^{\lambda}_{f}[\mu_{t}]\|_{L^{2}(\rho)}\leq R_{\lambda}$ and from the assumption on the regularity of $\phi$ it follows that:
$\|f^{*}(\Phi^{\top}\alpha)\|_{\mathcal{C}^{1,1}}\leq C.$
for some constant $C=C(f^{*},\lambda)$. Then by definition $\nabla\mathcal{L}^{\lambda}[\mu_{t}]=-\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu_{t}])\in\mathcal{C}^{0,1}$ and the first part of the result follows from classical results of ODE theory [hale2009ordinary] and on representation of solutions to continuity equation [ambrosio2008gradient, Thm. 8.1.8]. For the second part of the result, it suffices to remark that by the definition of ^$\hat{\mathcal{L}}^{\lambda}_{f}$ and $\nabla\mathcal{L}^{\lambda}_{f}$, for $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ and $j\in\{1,...,M\}$:
$M\nabla_{\omega_{j}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}\}_{1\leq i\leq M})=\nabla\mathcal{L}^{\lambda}_{f}[\hat{\mu}](\omega_{j})$ (35)
where $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. Therefore by eq. 28 we have that for a test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega)$:
$0=\frac{1}{M}\sum_{i=1}^{M}\int_{0}^{\infty}\frac{\mathrm{d}}{\mathrm{d}t}\varphi(t,\omega_{i}(t))\mathrm{d}t=\int_{0}^{\infty}\int_{\Omega}\left(\partial_{t}\varphi-\nabla\varphi\cdot\nabla\mathcal{L}^{\lambda}_{f}[\hat{\mu}_{t}]\right)\mathrm{d}\hat{\mu}_{t}\mathrm{d}t,$
meaning $(\hat{\mu}_{t})_{t\geq 0}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ according to Definition 4.1. ∎
Particle approximation
In the case where $\lambda>0$, associating the contraction rate of the gradient flow obtained in Theorem 4.1 with classical result on the approximation of measure by empirical distribution we obtain an approximation result for the minimization of $\mathcal{L}^{\lambda}_{f}$ with a finite number of features. For conciseness, we only state the result in the case $d\geq 3$, but similar results hold for $d\in\{1,2\}$.
Corollary 4.1 (Particle approximation).
Let the assumptions of Lemma 3.1 hold and let $d\geq 3$. Consider some initialization $\mu_{0}\in\mathcal{P}(\Omega)$ and, for some $N\geq 0$, note $\hat{\mu}_{0}\coloneqq N^{-1}\sum_{i=1}^{N}\delta_{\omega_{i}}$ the empirical measure where $\{\omega_{i}\}_{1\leq i\leq N}$ are i.i.d. samples of $\mu_{0}$. For $\lambda>0$, let $(\mu^{\lambda}_{t})_{t\geq 0}$ and $(\hat{\mu}_{t}^{\lambda})_{t\geq 0}$ be the gradient flow of $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ and $\hat{\mu}_{0}$ respectively. Then there exists a constant $A=A(d,\Omega)$ s.t. for every $t\geq 0$ and every $\varepsilon>0$:
$\mathbb{P}\left(\mathcal{W}_{1}(\hat{\mu}^{\lambda}_{t},\mu^{\lambda}_{t})\geq\varepsilon)\right)\leq\frac{A}{\varepsilon}N^{-1/d}e^{Ct}\,,$
where $C=C(f^{*},\lambda)$ is the constant in Theorem 4.1.
Proof.
Using [fournier2015rate, Thm. 1] we obtain at initialization $t=0$:
$\mathbb{E}\left[\mathcal{W}_{1}(\hat{\mu}^{\lambda}_{0},\mu^{\lambda}_{0})\right]\leq AN^{-1/d}$
for some constant $A=A(d,\Omega)>0$ depending on the dimension and on the domain $\Omega$. Then using the contraction rate in Theorem 4.1 we have a constant $C=C(f^{*},\lambda)>0$ such that for every $t\geq 0$:
$\mathbb{E}\left[\mathcal{W}_{1}(\hat{\mu}^{\lambda}_{t},\mu^{\lambda}_{t})\right]\leq AN^{-1/d}e^{Ct}.$
The result the follows by applying Markov’s inequality. ∎
4.2 Wasserstein gradient flows in the case $\lambda=0$ and ultra-fast diffusions
We now consider the limit of the proximal scheme eq. 30 when the step size $\tau$ tends to 00 and $\lambda$ is set to 00. We focus on the case where Assumption 1 holds and the regularization is of the form $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and recall that we use the shortcut $\mathcal{L}^{0}_{r}\coloneqq\mathcal{L}^{0}_{f}$. Then, following eq. 17, we have for $\mu\in\mathcal{P}(\Omega)$:
$\mathcal{L}^{0}_{r}(\mu)=\frac{1}{r-1}\mathrm{D}_{r}(\bar{\nu}|\mu)=\frac{1}{r-1}\int_{\Omega}\left|\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu=\frac{\|\bar{\nu}\|^{r}_{\mathrm{TV}}}{r-1}\int_{\Omega}\left|\frac{\mathrm{d}\bar{\mu}}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu.$ (36)
The first variation of $\mathcal{L}^{0}_{r}$ w.r.t. $\mu$ is formally given by $\frac{\delta\mathcal{L}^{0}_{r}}{\delta\mu}[\mu](\omega)=-\|\bar{\nu}\|^{r}_{\mathrm{TV}}\left(\frac{\bar{\mu}}{\mu}\right)^{r}$ and thus, following eq. 29, the Wasserstein gradient flow of $\mathcal{L}^{0}_{r}$ is formally defined as the solution to the continuity equation:
$\partial_{t}\mu_{t}=-\|\bar{\nu}\|^{r}\mathrm{div}\left(\mu_{t}\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r}\right).$ (37)
Moreover, calculating formally, $\nabla\left(\frac{\bar{\mu}}{\mu}\right)^{r}=r\frac{\bar{\mu}}{\mu}\nabla\left(\frac{\bar{\mu}}{\mu}\right)^{r-1}$ and eq. 37 can be written equivalently:
$\partial_{t}\mu_{t}=-r\|\bar{\nu}\|^{r}_{\mathrm{TV}}\mathrm{div}\left(\bar{\mu}\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r-1}\right)$
[…]
Definition 4.2 (Gradient flow of $\mathcal{L}^{0}_{r}$ (Def. 1.1 in [iacobelli2019weighted])).
Let $\mu_{0}\in\mathcal{P}(\Omega)$ admit a density $\mu_{0}\in L^{r+2}(\Omega)$. We say $(\mu_{t})_{t\geq 0}$ is a weak solution of eq. 37 or a gradient flow for $\mathcal{L}^{0}_{r}$ starting from $\mu_{0}$ if it a narrowly continuous curve in $\mathcal{P}(\Omega)$ with $\lim_{t\to 0^{+}}\mu_{t}=\mu_{0}$, s.t.
$\int_{0}^{\infty}\int_{\Omega}\left(\partial_{t}\varphi-\|\bar{\nu}\|^{r}_{\mathrm{TV}}\nabla\varphi\cdot\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r}\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega).$ (38)
and satisfying:
$\left(\frac{\mu_{t}}{\bar{\mu}}\right)^{r-1}\in L^{2}_{loc}((0,\infty),H^{1}(\Omega)),\quad\frac{\bar{\mu}}{\mu_{t}}\in L^{2}_{loc}((0,\infty),H^{1}(\Omega))\,.$
Existence and uniqueness of solutions
In [iacobelli2019weighted], the authors establish the existence of uniqueness of gradient flows for the functional $\mathcal{L}^{0}_{r}$. More precisely, they show that, under appropriate assumptions on the initialization $\mu_{0}$ and on the target $\bar{\mu}$, the iterates of the proximal scheme in eq. 30 converge towards a curve $(\mu_{t})_{t\geq 0}$ that is a gradient flow of the functional $\mathcal{L}^{0}_{r}$ in the sense of Definition 4.2.
Theorem 4.2 (Well-posedness of the gradient flow equation for $\lambda=0$ (Thm. 1.2 of [iacobelli2019weighted])).
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous and have bounded log-densities. Then there exists a unique weak solution of eq. 37 starting from $\mu_{0}$ in the sense of Definition 4.2.
Convergence towards the target distribution
[…]
Theorem 4.3 (Theorem 1.4 in [iacobelli2019weighted]).
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous and have bounded log-densities. For $\mu_{0}\in\mathcal{P}(\Omega)$, let $(\mu_{t})_{t\geq 0}$ be a weak solution of eq. 37 starting from $\mu_{0}$ in the sense of Definition 4.2. Then in fact the log-density of $\mu_{t}$ is bounded uniformly over $t\geq 0$ and there exists a constant $C=C(\Omega,\bar{\mu},\mu_{0})>0$ s.t. for every $t\geq 0$:
$\|\bar{\mu}-\mu_{t}\|_{L^{2}(\Omega)}\leq Ce^{-Ct}.$
For completeness, we give here some of the key arguments of the proof of the above Theorem 4.3 in the case where $r=2$. In this case we have for every $\mu\in\mathcal{P}(\Omega)$:
$\mathcal{L}^{0}_{2}(\mu)=\|\bar{\nu}\|^{2}_{\mathrm{TV}}\int_{\Omega}\left|\frac{\mathrm{d}\bar{\mu}}{\mathrm{d}\mu}\right|^{2}\mathrm{d}\mu=\|\bar{\nu}\|^{2}_{\mathrm{TV}}\left(\chi^{2}(\bar{\mu}|\mu)+1\right),$ (39)
where $\chi^{2}$ is the chi-square divergence. The following Lemma 4.3 establishes the desired linear convergence rate for the proximal scheme defined in eq. 30 with the loss $\mathcal{L}^{0}_{2}$. The result in continuous time then follows then follows from the lower semi-continuity of the $\chi^{2}$-divergence as the curve $(\mu_{t})_{t\geq 0}$ is obtained by taking the limit of the discrete process $(\mu^{\tau}_{k})_{k\geq 0}$ when the discretization time $\tau$ tends to zero.
From a technical perspective, the proof of Lemma 4.3 relies on a Poincaré inequality satisfied by $\mu_{t}$. It is indeed well-known that such inequality controls the convergence rate of Fokker-Planck equations towards their stationary distribution distribution in $\chi^{2}$-distance [pavliotis2014stochastic, Thm. 4.4]. This can for example be used to prove the convergence of sampling algorithms such as Langevin Monte Carlo [chewi2024analysis, chewi2020svgd]. In our case, the ultra-fast diffusion eq. 37 is to be interpreted as a Wasserstein gradient flow for $\mathcal{L}^{0}_{2}$, which, by the above eq. 39, is the reverse $\chi^{2}$-divergence between $\mu_{t}$ and $\bar{\mu}$ and the convergence rate is controlled by the Poincaré constant of $\mu_{t}$. This rate may thus a priori evolve and vanish during training but, crucially, [iacobelli2019weighted, Lem. 2.4] shows that it is here a property of solutions to the ultra-fast diffusion equation that the log-density ratio $\|\log\left(\frac{\bar{\mu}}{\mu_{t}}\right)\|_{\infty}$ decreases with time. As a consequence, it is sufficient to assume that the log-density is bounded at initialization to obtain a control over the Poincaré constant of $\mu_{t}$, for $t\geq 0$, by a classical perturbation argument [ane2000inegalites, Thm. 3.4.1].
Lemma 4.3.
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous with bounded log-densities. Let $\tau>0$ and let $(\mu_{k}^{\tau})_{k\geq 0}$ be the sequence defined by eq. 30 with $\lambda=0$, $f(t)=|t|^{2}$ ($r=2$) and initialization $\mu_{0}^{\tau}=\mu_{0}\in\mathcal{P}(\Omega)$. Then there exists a constant $C>0$ s.t.:
$\forall k\geq 0,\quad\chi^{2}(\bar{\mu}|\mu^{\tau}_{k})\leq(1+C\tau)^{-k}\chi^{2}(\bar{\mu}|\mu_{0}).$
Proof.
From [iacobelli2019weighted, Thm.2.1 and Lem.2.4] we know the sequence $(\mu^{\tau}_{k})_{k\geq 0}$ is uniquely defined. Moreover $\mu^{\tau}_{k}$ is absolutely continuous w.r.t. Lebesgue measure and their exists a constant $C=C(\bar{\mu},\mu_{0})>0$ s.t. the $\log$-densities $\log(\mu^{\tau}_{k})$ satisfy:
$\forall k\geq 0,\quad\left\|\log(\mu^{\tau}_{k})\right\|_{\infty}\leq C.$
Then, at step $k\geq 0$, we get from the expression of $\mathcal{L}^{0}_{2}$ in eq. 39 and from the optimality condition in eq. 30 (see e.g. [santambrogio2015optimal, Prop.7.20]) that:
$-\|\bar{\nu}\|^{2}_{\mathrm{TV}}\left(\frac{\bar{\mu}}{\mu^{\tau}_{k+1}}\right)^{2}+\frac{\varphi}{\tau}=cte,\quad\text{almost everywhere on $\Omega$,}$
where $\varphi$ is the Kantorovitch potential from $\mu^{\tau}_{k+1}$ to $\mu^{\tau}_{k}$. Also this potential is necessarily Lipschitz, hence a.e. differentiable and so is $\bar{\nu}/\mu^{\tau}_{k+1}$. Then from the definition of $\mu_{k+1}^{0,\tau}$ we have:
$\mathcal{L}^{0}_{2}(\mu^{\tau}_{k})-\mathcal{L}^{0}_{2}(\mu^{\tau}_{k+1})$ $\geq\frac{1}{2\tau}\mathcal{W}_{2}(\mu^{\tau}_{k+1},\mu^{\tau}_{k})^{2}$
$=\frac{1}{2\tau}\int_{\Omega}\left|\nabla\varphi\right|^{2}\mathrm{d}\mu^{\tau}_{k+1}$
$=\frac{\tau\|\bar{\nu}\|^{2}_{\mathrm{TV}}}{2}\int_{\Omega}\left\|\nabla\left(\frac{\bar{\mu}}{\mu^{\tau}_{k+1}}\right)^{2}\right\|^{2}\mathrm{d}\mu^{\tau}_{k+1}$
where we used the definition of the potential $\varphi$ and the optimality condition. Using that $\mu^{\tau}_{k+1}$ has log-density bounded by $C=C(\bar{\mu},\mu_{0})$ and that the domain $\Omega$ satisfies a Poincaré inequality with constant $C_{P}=C_{P}(\Omega)$, it follows from a classical perturbation argument that $\mu^{\tau}_{k+1}$ satisfies a Poincaré inequality with constant $e^{2C}C_{P}(\Omega)$ [ane2000inegalites, Thm. 3.4.1]. As a consequence:
$\int_{\Omega}\left\|\nabla\left(\frac{\bar{\mu}}{\mu^{\tau}_{k+1}}\right)^{2}\right\|^{2}\mathrm{d}\mu^{\tau}_{k+1}$ $\geq 4e^{-4C}\int_{\Omega}\left\|\nabla\left(\frac{\bar{\mu}}{\mu^{\tau}_{k+1}}\right)\right\|^{2}\mathrm{d}\mu^{\tau}_{k+1}$
$\geq 4C_{P}^{-1}e^{-6C}\left(\int_{\Omega}\left(\frac{\bar{\mu}}{\mu^{\tau}_{k+1}}\right)^{2}\mathrm{d}\mu^{\tau}_{k+1}-1\right)$
$=4C_{P}^{-1}e^{-6C}\|\bar{\nu}\|^{-2}_{\mathrm{TV}}\left(\mathcal{L}^{0}_{2}(\mu^{\tau}_{k+1})-\|\bar{\nu}\|^{2}_{\mathrm{TV}}\right),$
where $\|\bar{\nu}\|^{2}_{\mathrm{TV}}=\inf\mathcal{L}^{0}_{2}$. Combining this with the previous inequality finally gives:
$\left(1+2\tau C_{P}^{-1}e^{-6C}\right)\left(\mathcal{L}^{0}_{2}(\mu^{\tau}_{k+1})-\inf\mathcal{L}^{0}_{2}\right)\leq\mathcal{L}^{0}_{2}(\mu^{\tau}_{k})-\inf\mathcal{L}^{0}_{2},$
and inductively:
$\forall k\geq 0,\quad\mathcal{L}^{0}_{2}(\mu^{\tau}_{k})-\inf\mathcal{L}^{0}_{2}\leq\left(1+2\tau C_{P}^{-1}e^{-6C}\right)^{-k}\left(\mathcal{L}^{0}_{2}(\mu_{0})-\inf\mathcal{L}^{0}_{2}\right).$
By the definition of $\mathcal{L}^{0}_{2}$ in eq. 39, this is the desired result. ∎
Remark 4.2 (Dependence of the convergence rate w.r.t. dimension).
It follows from the proof that the convergence rate $C$ in Lemma 4.3 scales linearly with $C_{P}(\Omega)^{-1}$ where $C_{P}(\Omega)$ is the Poincaré constant of the domain $\Omega$. For bounded, Lipschitz and convex domains of $\mathbb{R}^{d}$ or for the flat Torus $\mathbb{T}^{d}$, this constant is in particular independent of the dimension $d$ [payne1960optimal]. Therefore, the correspondence between the training of neural networks in the two-timescale regime and solutions to ultra-fast diffusions points towards the fact that gradient methods, with suitable hyperparameter scaling, are amenable to efficient feature learning in the training of neural networks, without suffering from the curse of dimensionality [donoho2000high]. Note however that the convergence rate $C$ in Lemma 4.3 is exponentially bad in the log-density ratio $\|\log(\bar{\mu}/\mu_{0})\|_{\infty}$. In particular the convergence rate does not hold in case the teacher feature distribution is supported on a finite number of atoms.
5 Convergence of gradient flow
The main purpose of this work is to study in what extend the gradient flow dynamics defined Definitions 4.1 and 4.2 allow to recover the teacher feature distribution $\bar{\mu}$ associated to the observed signal $Y$ in Assumption 1. Whereas Theorem 4.3 shows convergence of the gradient flow of $\mathcal{L}^{0}_{f}$, that is solutions to the ultra-fast diffusion eq. 37, such dynamics are hardly numerically tractable in practice due to the absence of the regularization parameter $\lambda$. For this reason we are interested here in the asymptotic behavior of the gradient flow of $\mathcal{L}^{\lambda}_{f}$ in the case where $\lambda>0$. A difficulty is that, in the case $\lambda=0$, the proof of Theorem 4.3 relies on the implicit behavior of eq. 37 which preserves the density of solutions. Such a behavior is a priori not expected to hold when $\lambda>0$. As a consequence, the result of this section holds under supplementary regularity assumptions on the solutions to eq. 33.
5.1 Algebraic convergence rate
At fixed $\lambda>0$, we are able to obtain convergence towards the minimizer $\bar{\mu}^{\lambda}_{f}$ of $\mathcal{L}^{\lambda}_{f}$ under mild regularity assumptions on solutions to the gradient flow eq. 33. Specifically, for a probability measure $\mu\in\mathcal{P}(\Omega)$ and a function $h\in\mathcal{C}^{1}$ we define the weighted Sobolev semi-norm of $h$ as:
$\|h\|_{\dot{H}^{1}(\mu)}\coloneqq\left(\int_{\Omega}\|\nabla h\|^{2}\mathrm{d}\mu\right)^{1/2}.$
Then for a measure $\nu\in\mathcal{M}(\Omega)$ s.t. $\int_{\Omega}\mathrm{d}\nu=0$ the negative weighted Sobolev semi-norm $\|\nu\|_{\dot{H}^{-1}(\mu)}$ is defined by duality as:
$\|\nu\|_{\dot{H}^{-1}(\mu)}\coloneqq\sup_{\|h\|_{\dot{H}^{1}(\mu)}\leq 1}\int_{\Omega}h\mathrm{d}\nu.$
The following Theorem 5.1 states convergence of $\mathcal{L}^{\lambda}_{f}$ towards 00 with an algebraic convergence rate provided $\|\mu_{t}-\frac{\bar{\nu}}{\bar{m}}\|_{\dot{H}^{-1}(\mu_{t})}$ stays bounded along the gradient flow. As discussed below, since the domain $\Omega$ is compact, this assumption is satisfied for example when both distribution have bounded log-densities. The arguments are similar to the one presented in [glaser2021kale], where the authors consider an infimal convolution between a kernel discrepancy and the $\mathrm{KL}$-divergence. Importantly, the obtained convergence rate depends on the bound on $\|\mu_{t}-\frac{\bar{\nu}}{\bar{m}}\|_{\dot{H}^{-1}(\mu_{t})}$ but is independent of $\lambda>0$.
Theorem 5.1.
Let Assumption 1 hold. Consider $\lambda>0$ and some initialization $\mu_{0}\in\mathcal{P}(\Omega)$. Let $(\mu_{t})_{t\geq 0}$ be the gradient flow of $\mathcal{L}_{f}^{\lambda}$, starting from $\mu_{0}$ (in the sense of eq. 33). Assume that:
  • $\bar{\nu}$ is a positive measure and $f$ is such that $\min f=f(\bar{m})=0$ where $\bar{m}\coloneqq\bar{\nu}(\Omega)>0$.
  • the gradient flow $(\mu_{t})_{t\geq 0}$ is such that $\|\frac{\bar{\nu}}{\bar{m}}-\mu_{t}\|_{\dot{H}^{-1}(\mu_{t})}$ is bounded, uniformly over $t\geq 0$.
Then there exists a constant $C>0$ s.t. for any $t\geq 0$:
$\mathcal{L}^{\lambda}_{f}(\mu_{t})\leq\left(\mathcal{L}^{\lambda}_{f}(\mu_{0})^{-1}+Ct\right)^{-1}.$
In particular, $\mu_{t}$ converges to $\bar{\mu}=\bar{\nu}/\bar{m}$ when $t$ tends to +$\infty$+ ∞.
Proof.
Note that it follows from the assumptions on $f$ and $\bar{\nu}$ that $\inf\mathcal{L}^{\lambda}_{f}=0$ and that this infimum is attained only for $\mu=\bar{\nu}/\bar{m}$. Thus, the last statement on the convergence of $\mu_{t}$ follows from the convergence of $\mathcal{L}^{\lambda}_{f}(\mu_{t})$ to 00 and from the lower-semicontinuity of $\mathcal{L}^{\lambda}_{f}$ (see section 3).
To obtain the convergence rate, consider $\mu\in\mathcal{P}(\Omega)$ and note that by eq. 19 we have $\mathcal{L}^{\lambda}_{f}(\mu)=\max_{\alpha}\mathcal{K}(\alpha,\mu)$ where for every $\alpha\in L^{2}(\rho)$:
$\mathcal{K}(\alpha,\mu)\coloneqq\int_{\Omega}(\Phi^{\top}\alpha)\mathrm{d}\bar{\nu}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)},$
where $f^{*}$ is the Legendre transform of $f$. Let us note $\alpha^{\lambda}=\alpha^{\lambda}_{f}[\mu]$ the maximizer of $\mathcal{K}(\alpha,\mu)$. Then, using the convexity of $f^{*}$ we have for every $\omega\in\Omega$:
$f^{*}(0)+\partial f^{*}(0)(\Phi^{\top}\alpha^{\lambda})(\omega)\leq f^{*}((\Phi^{\top}\alpha^{\lambda})(\omega)).$
Also by assumptions $\partial f(\bar{m})=0$ and hence by properties of the Legendre transform $\partial f^{*}(0)=\bar{m}$. Also $f^{*}(0)=-f(\bar{m})=0$ and after integrating w.r.t. $\bar{\nu}$ :
$\int_{\Omega}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\bar{\nu}\leq\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\frac{\mathrm{d}\bar{\nu}}{\bar{m}}.$
Then, replacing $\alpha$ by $\alpha^{\lambda}$ in $\mathcal{K}$ and using the previous inequality:
$\mathcal{L}^{\lambda}_{f}(\mu)=\mathcal{K}(\alpha^{\lambda},\mu)\leq\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}(\frac{\bar{\nu}}{\bar{m}}-\mu)\leq\left\|f^{*}(\Phi^{\top}\alpha^{\lambda})\right\|_{\dot{H}^{1}(\mu)}\|\frac{\bar{\nu}}{\bar{m}}-\mu\|_{\dot{H}^{-1}(\mu)}$
Also, by the gradient flow equation eq. 33, the dissipation of $\mathcal{L}^{\lambda}_{f}$ along the gradient flow curve $(\mu_{t})_{t\geq 0}$ is given for every $t\geq 0$ by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{L}^{\lambda}_{f}(\mu_{t})=-\int_{\Omega}\left\|\nabla(f^{*}(\Phi^{\top}\alpha^{\lambda}_{t}))\right\|^{2}\mathrm{d}\mu_{t}=-\left\|f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})\right\|^{2}_{\dot{H}^{1}(\mu_{t})},$
where $\alpha^{\lambda}_{t}=\alpha^{\lambda}_{f}[\mu_{t}]$ maximizes $\mathcal{K}(\alpha,\mu_{t})$. Thus using the previous inequality on $\mathcal{L}^{\lambda}_{f}(\mu_{t})$ and that $\|\frac{\bar{\nu}}{\bar{m}}-\mu_{t}\|_{\dot{H}^{-1}(\mu_{t})}$ is bounded, uniformly over $t\geq 0$, we get for every $t\geq 0$:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{L}^{\lambda}_{f}(\mu_{t})\leq-C^{-1}\mathcal{L}^{\lambda}_{f}(\mu_{t})^{2},$
for some constant $C>0$. The desired convergence rate follows from this inequality by applying a Grönwall lemma. ∎
Let us comment on the assumptions of Theorem 5.1. The second assumption specifically, is automatically satisfied in case $\bar{\nu}$ has bounded density and $\mu_{t}$ has bounded log-density, uniformly over $t\geq 0$. Indeed, for $\mu\in\mathcal{P}(\Omega)$ having a lower-bounded log-density we have that the weighted Sobolev semi-norm .$\|.\|_{\dot{H}^{1}(\mu)}$ is lower-bounded by the classical Sobolev semi-norm .$\|.\|_{\dot{H}^{1}(\pi)}$ where we recall that $\pi$ is the (normalized) Lebesgue measure over $\Omega$. Precisely, if $\pi\ll\mu$ and $\mathrm{d}\pi/\mathrm{d}\mu\leq C_{1}$ then for every $f\in\mathcal{C}^{1}$:
$\|f\|_{\dot{H}^{1}(\pi)}\leq C_{1}\|f\|_{\dot{H}^{1}(\mu)}.$
In this case, the weighted negative Sobolev semi-norm .$\|.\|_{\dot{H}^{-1}(\mu)}$ is upper-bounded by the semi-norm .$\|.\|_{\dot{H}^{-1}(\pi)}$ and for every $\nu\in\mathcal{M}(\Omega)$ with $\int_{\Omega}\mathrm{d}\nu=0$ we have:
$\|\nu\|_{\dot{H}^{-1}(\mu)}\leq C_{1}\|\nu\|_{\dot{H}^{-1}(\pi)}.$
Moreover, this last quantity can be estimated by the Wasserstein distance. Indeed, for probability measures having bounded log-densities, the Wasserstein distance $\mathcal{W}_{2}$ is equivalent to the negative Sobolev semi-norm .$\|.\|_{\dot{H}^{-1}(\pi)}$. If $\mu,\nu\in\mathcal{P}(\Omega)$ are such that $\frac{\mathrm{d}\mu}{\mathrm{d}\pi},\frac{\mathrm{d}\nu}{\mathrm{d}\pi}\leq C_{2}$ for some constant $C_{2}>0$ we have [santambrogio2015optimal, Lem. 5.33 and Thm. 5.34]:
$\|\mu-\nu\|_{\dot{H}^{-1}(\pi)}\leq C_{2}^{1/2}\mathcal{W}_{2}(\mu,\nu).$
Finally, the Wasserstein distance $\mathcal{W}_{2}(\mu,\nu)$ is always bounded by $\mathrm{diam}(\Omega)$ which is finite, hence ensuring the second assumption of Theorem 5.1 is satisfied.
5.2 Convergence to ultra-fast diffusion.
The algebraic convergence rate stated in the above Theorem 5.1 in the case $\lambda>0$ stands in contrast with the faster linear convergence stated in Theorem 4.3 in the case $\lambda=0$. For this reason, we are interested in comparing the gradient flow dynamics with and without regularization.
Below we assume $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and Theorem 5.2 shows local uniform in time convergence of gradient flows of $\mathcal{L}^{\lambda}_{r}$ to gradient flows of $\mathcal{L}^{0}_{r}$, i.e. solutions to the ultra-fast diffusion equation eq. 36, when the regularization strength $\lambda$ vanishes. To obtain such a result we assume regularity on the density ratio $\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}}$. Namely, we assume that the Legendre-conjugate $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})$ stays bounded in the RKHS $\mathcal{H}$, defined as the image of the convolution operator $\Phi^{\top}:L^{2}(\rho)\to\mathcal{C}^{0}(\Omega)$ (eq. 41). Using classical results from the theory of inverse problems, such a source condition ensures the dual variable $\alpha\in L^{2}(\rho)$ stays uniformly bounded for $\lambda>0$ (Lemma 5.1). Therefore, provided $\mathcal{H}$ is sufficiently regular, such a regularity assumption ensures compactness of the Wasserstein gradient $\nabla\mathcal{L}^{\lambda}_{r}[\mu_{t}]=\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})$ in $\mathcal{C}^{1}$ and allows passing to the limit in eq. 34 to obtain eq. 38.
Theorem 5.2.
Assume Assumption 1 hold with $\bar{\nu}$ a positive measure with bounded log-density, $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and the assumptions of Theorem 4.1 and Theorem 4.2 are satisfied. Consider some initialization $\mu_{0}\in\mathcal{P}(\Omega)$ s.t. $\mu_{0}$ has bounded log-density. For $\lambda\geq 0$ let $(\mu^{\lambda}_{t})_{t\geq 0}$ be the gradient flow of $\mathcal{L}^{\lambda}_{r}$, starting from $\mu_{0}$ in the sense of Definition 4.1 (when $\lambda>0$) and Definition 4.2 (when $\lambda=0$). Moreover, assume $\mathcal{H}$ is compactly embedded in $\mathcal{C}^{1}(\Omega)$ and $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})$ is bounded in $\mathcal{H}$, locally uniformly over $t\geq 0$ and uniformly over $\lambda>0$. Then for any $T\geq 0$:
$\lim_{\lambda\to 0^{+}}\sup_{t\in[0,T]}\mathcal{W}_{2}(\mu_{t}^{0},\mu_{t}^{\lambda})=0.$
Proof.
The curve $(\mu_{t})^{\lambda}$ for $\lambda>0$ are gradient flows for the functionals $\mathcal{L}^{\lambda}_{r}$ and classical computations shows that for every $\lambda\geq 0$ and $t,s\geq 0$:
$\mathcal{W}_{2}(\mu^{\lambda}_{t},\mu^{\lambda}_{s})^{2}\leq|t-s|\left|\mathcal{L}^{\lambda}_{r}(\mu^{\lambda}_{t})-\mathcal{L}^{\lambda}_{r}(\mu^{\lambda}_{s})\right|\leq|t-s|\mathcal{L}^{0}_{r}(\mu_{0}),$
where we used the fact that the functionals $\mathcal{L}^{\lambda}_{r}$ converge pointwise from below to $\mathcal{L}^{0}_{r}$. Thus, for $T\geq 0$, the sequence $(\mu_{t}^{\lambda})_{t\in[0,T]}$ is uniformly equicontinuous with value in the compact space $\mathcal{P}(\Omega)$ and Arzela-Ascoli’s theorem ensures the existence of a subsequence $\lambda_{n}\to 0^{+}$ s.t.:
$(\mu_{t}^{\lambda})_{t\in[0,T]}\xrightarrow{n\to\infty}(\mu_{t})_{t\in[0,T]}\in\mathcal{C}^{0}([0,T],\mathcal{P}(\Omega)).$
To prove the result one needs to identify $\mu_{t}$ with $\mu_{t}^{0}$ and the supplementary regularity assumptions on $\mu_{t}^{\lambda}$ are sufficient for this purpose. Let us fix some $t\in[0,T]$ and note $u^{\lambda}_{t}=u^{\lambda}_{f}[\mu^{\lambda}_{t}]$ the minimizer in eq. 9, $\nu^{\lambda}_{t}\in\mathcal{M}(\Omega)$ the minimizer in eq. 18 s.t. $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}=u^{\lambda}_{t}$ and $\alpha^{\lambda}_{t}\in L^{2}(\rho)$ the maximizer in eq. 19.
Then for every $\lambda>0$, since $\mu_{0}$ has bounded log-density we have by the flow-map representation in Proposition 4.1 that $\mu^{\lambda}_{t}$ has bounded log-density. Also, since $\bar{\nu}$ is positive with bounded log-density and $\Phi\star$ is injective, we have that $u^{\dagger}_{t}\coloneqq\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}}$ is the unique solution to eq. 10. But then, by definition of the RKHS $\mathcal{H}$ we have that $\Phi^{\top}:L^{2}(\rho)\to\mathcal{H}$ is a partial isometry by Mercer’s theorem (Theorem A.1) and the assumption that $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})\in\mathcal{H}$ is equivalent to a source condition of the form in eq. 40. Hence by Lemma 5.1, the dual variable $\alpha^{\lambda}_{t}$ is bounded in $L^{2}(\rho)$, uniformly over $\lambda>0$, which implies that (up to extraction of a subsequence) $\Phi^{\top}\alpha^{\lambda}_{t}$ converges to some $h_{t}$ in $\mathcal{C}^{1}(\Omega)$.
Also for every $\lambda>0$, by the duality relations in eq. 20, we have $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}=\partial f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})$ and hence (recall $f(t)=|t|^{r}/(r-1)$ for some $r>1$) $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}\to\partial f^{*}(h_{t})$ in $\mathcal{C}^{0}(\Omega)$. Since $\mathcal{L}^{\lambda}_{r}(\mu^{\lambda}_{t})\leq\mathcal{L}^{0}_{r}(\mu_{0})$ is bounded we have by eq. 18 that $\nu^{\lambda}_{t}\to\bar{\nu}$ narrowly and then for every $\varphi\in\mathcal{C}^{0}(\Omega)$:
$\int_{\Omega}\varphi\mathrm{d}\nu^{\lambda}_{t}=\int_{\Omega}\varphi\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}\mathrm{d}\mu^{\lambda}_{t}\xrightarrow{\lambda\to 0^{+}}\int_{\Omega}\varphi\mathrm{d}\bar{\nu}=\int_{\Omega}\varphi\partial f^{*}(h_{t})\mathrm{d}\mu_{t}.$
This shows that $\bar{\nu}$ is absolutely continuous w.r.t. $\mu_{t}$ and that $\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu_{t}}=\partial f^{*}(h_{t})$. By duality, this is equivalent to $h_{t}=\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu_{t}})$, which shows that $\Phi^{\top}\alpha^{\lambda}_{t}$ converges to $h_{t}$ in $\mathcal{C}^{1}(\Omega)$.
Finally, using the gradient flow equation in eq. 34, the previously described convergence of $\Phi^{\top}\alpha^{\lambda}_{t}$ is sufficient to have for every test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,T)\times\Omega)$:
$\int_{0}^{T}\int_{\Omega}\left(\partial_{t}\varphi_{t}-\frac{1}{2}\nabla\varphi_{t}\cdot\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})\right)\mathrm{d}\mu^{\lambda}_{t}\mathrm{d}t=0$
$\xrightarrow{\lambda\to 0^{+}}$ $\int_{0}^{T}\int_{\Omega}\left(\partial_{t}\varphi_{t}-\frac{1}{2}\nabla\varphi_{t}\cdot\nabla f^{*}(\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu_{t}}))\right)\mathrm{d}\mu_{t}\mathrm{d}t=0.$
Since $f(t)=|t|^{r}/(r-1)$ for some $r>1$ the above equation is equivalent to eq. 38 which shows $\mu_{t}$ is the weak solution starting from $\mu_{0}$ of the ultra-fast diffusion equation eq. 36 according to Definition 4.2, that is $\mu_{t}=\mu^{0}_{t}$. ∎
The proof of the above Theorem 5.2 relies on the following result on solutions to inverse problems with nonlinear regularization [benning2018modern]. The following Lemma 5.1 is similar to [iglesias2018note, Prop. 3].
Lemma 5.1.
Assume $f$ satisfies Assumption 2 and Assumption 1 holds. For $\mu\in\mathcal{P}(\Omega)$, let $u^{\dagger}\in L^{1}(\mu)$ be a solution of eq. 10. We say $u^{\dagger}$ satisfies a source condition if there exists $\alpha\in L^{2}(\rho)$ s.t.
$\Phi^{\top}\alpha\in\partial f(u^{\dagger})\quad\text{in $L^{1}(\mu)$}.$ (40)
Then in this case, noting $\alpha^{\dagger}\in L^{2}(\mu)$ the $\alpha$ of minimal norm satisfying the above source condition, we have for every $\lambda>0$
$\|\alpha^{\lambda}_{f}[\mu]\|_{L^{2}(\rho)}\leq\|\alpha^{\dagger}\|_{L^{2}(\rho)}\quad\text{and}\quad\alpha^{\lambda}_{f}[\mu]\xrightarrow{\lambda\to 0^{+}}\alpha^{\dagger}$
where $\alpha^{\lambda}_{f}[\mu]$ is the solution to eq. 19.
Proof.
Let $u^{\dagger}\in L^{1}(\mu)$ and $\alpha^{\dagger}\in L^{2}(\rho)$ be as in the statement. By the source condition eq. 40 we have in $L^{1}(\mu)$:
$-f^{*}(\Phi^{\top}\alpha^{\dagger})+(\Phi^{\top}\alpha^{\dagger})u^{\dagger}\geq f(u^{\dagger})$
and integrating w.r.t. $\mu$ and using that $\int_{\Omega}(\Phi^{\top}\alpha^{\dagger})u^{\dagger}\mathrm{d}\mu=\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}$ we obtain:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}\geq\int_{\Omega}f(u^{\dagger})\mathrm{d}\mu=\inf_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu.$
Thus, $\alpha^{\dagger}$ achieves the supremum in eq. 19 with $\lambda=0$ and we have for every other $\alpha\in L^{2}(\rho)$:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)}.$
Moreover, for $\lambda>0$, noting $\alpha^{\lambda}\coloneqq\alpha^{\lambda}_{f}[\mu]$, we have by definition:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\mu+\langle\alpha^{\lambda},Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\lambda}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\dagger}\|^{2}_{L^{2}(\rho)}.$
Substracting the two previous inequalities and simplifying gives:
$\|\alpha^{\lambda}\|_{L^{2}(\rho)}\leq\|\alpha^{\dagger}\|_{L^{2}(\rho)}.$
Thus $\alpha^{\lambda}$ is bounded, uniformly over $\lambda>0$. For the convergence part, note that since it is bounded in $L^{2}(\rho)$ it converges weakly to some $\alpha^{0}\in L^{2}(\rho)$. Also taking the optimality condition for $\alpha^{\lambda}$ and taking the limit when $\lambda\to 0^{+}$ we obtain for every $\alpha\in L^{2}(\rho)$:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\mu+\langle\alpha^{\lambda},Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\lambda}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)}$
$\xrightarrow{\lambda\to 0+}$ $-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{0})\mathrm{d}\mu+\langle\alpha^{0},Y\rangle_{L^{2}(\rho)}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)},$
which shows $\alpha^{0}$ is also a maximizer of the dual problem eq. 19 when $\lambda=0$ and, as a consequence also satisfies the source condition eq. 40. But, by minimality of the norm of $\alpha^{\dagger}$ and by weak lower-semicontinuity of the norm we have:
$\|\alpha^{\dagger}\|_{L^{2}(\rho)}\leq\|\alpha^{0}\|_{L^{2}(\rho)}\leq\liminf_{\lambda\to 0^{+}}\|\alpha^{\lambda}\|_{L^{2}(\rho)}\leq\|\alpha^{\dagger}\|_{L^{2}(\rho)}$
which shows that in fact $\alpha^{\lambda}\to\alpha^{\dagger}$ strongly in $L^{2}(\rho)$. ∎
Acknowledgements
The work of G. Peyré was supported by the European Research Council (ERC project WOLF) and the French government under the management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PRAIRIE-PSAI). The work of F.-X. Vialard was supported by the Bézout Labex (New Monge Problems), funded by ANR, reference ANR-10-LABX-58.
[…]
Appendix A Positive definite kernels and RKHS
We recall in this section basic properties on the theory of Reproducing Kernel Hilbert Spaces and refer to classical textbooks for a complete presentation of the topic [steinwart2008support, scholkopf2002learning]. In this work we consider a mapping $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$ as well as a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$. This choice of $\phi$ and $\rho$ determines a symmetric, positive definite kernel [steinwart2008support, Def. 4.15] $\kappa:\Omega\times\Omega\to\mathbb{R}$ defined by:
$\forall\omega,\omega^{\prime}\in\Omega,\quad\kappa(\omega,\omega^{\prime})\coloneqq\int_{X}\phi(\omega,x)\phi(\omega^{\prime},x)\mathrm{d}\rho(x)=\langle\phi(\omega,.),\phi(\omega^{\prime},.)\rangle_{L^{2}(\rho)}.$
Thus, associated to $\kappa$ is a (uniquely defined) structure of Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$ with scalar product $\langle.,.\rangle_{\mathcal{H}}$, that is a Hilbert space of functions on $\Omega$ such that [steinwart2008support, Def. 4.18]: (i) $\kappa(\omega,.)\in\mathcal{H}$ for every $\omega\in\Omega$ and (ii) the following reproducing property holds:
$\forall h\in\mathcal{H},\omega\in\Omega,\quad h(\omega)=\langle h,\kappa(\omega,.)\rangle_{\mathcal{H}}.$
Following the definition of $\kappa$, $L^{2}(\rho)$ is a feature space for $\mathcal{H}$ and $\phi$ a feature map [steinwart2008support, Def. 4.1]. Also, $\mathcal{H}$ can be isometrically identified as a subspace of $L^{2}(\rho)$ and the convolution with $\phi$ is a partial isometry [steinwart2008support, Thm. 4.21]. Precisely, we have =$\mathcal{H}=\{h:\Omega\to\mathbb{R}\;|\;\exists\alpha\in L^{2}(\rho),h=\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x)\}$ and the RKHS norm on $\mathcal{H}$ satisfies:
$\forall h\in\mathcal{H},\quad\|h\|_{\mathcal{H}}=\inf\left\{\|\alpha\|_{L^{2}(\rho)}\;|\;h=\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x)\right\}.$ (41)
In this paper, we always work with the following minimal assumption on the feature map $\phi$:
Assumption 3 (Assumption on $\phi$).
The feature map $\phi$ is in $L^{2}(\rho,\mathcal{C}^{0})$. In particular, it implies the kernel $\kappa$ is continuous on $\Omega\times\Omega$.
Kernel embeddings and kernel discrepancy between measures
The above assumption is sufficient to ensure $\mathcal{H}$ is included in $\mathcal{C}(\Omega)$ and guarantees the existence of kernel embeddings for finite Borel measures [muandet2017kernel, gretton2012kernel]. For a measure $\nu\in\mathcal{M}(\Omega)$ its kernel embedding $M_{\kappa}(\nu)$ is defined as the unique element of $\mathcal{H}$ satisfying:
$\forall h\in\mathcal{H},\quad\int_{\Omega}h\mathrm{d}\nu=\langle h,M_{\kappa}(\nu)\rangle_{\mathcal{H}}.$ (42)
Equivalently, the kernel embedding is given by the Böchner intgral $M_{\kappa}(\nu)=\int_{\Omega}\kappa(.,\omega)\mathrm{d}\nu(\omega)\in\mathcal{H}$. This embedding defines a discrepancy between measures by seeing them as element of the Hilbert space $\mathcal{H}$. For two measures $\nu,\nu^{\prime}\in\mathcal{M}(\Omega)$ the Maximum Mean Discrepancy (MMD) between $\nu$ and $\nu^{\prime}$ is defined as [muandet2017kernel, gretton2012kernel]:
$\mathrm{MMD}_{\kappa}(\nu,\nu^{\prime})=\left\|M_{\kappa}(\nu)-M_{\kappa}(\nu^{\prime})\right\|_{\mathcal{H}}.$
Alternatively, Assumption 3 is sufficient to ensure the convolution $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ defined in eq. 7 is a bounded operator and by construction we have:
$\mathrm{MMD}_{\kappa}(\nu,\nu^{\prime})=\left(\iint_{\Omega\times\Omega}\kappa(\omega,\omega^{\prime})\mathrm{d}(\nu-\nu^{\prime})(\omega)\mathrm{d}(\nu-\nu^{\prime})(\omega^{\prime})\right)^{1/2}=\left\|\Phi\star(\nu-\nu^{\prime})\right\|_{L^{2}(\rho)}.$ (43)
The discrepancy $\mathrm{MMD}_{\kappa}$ is in particular a distance between measures whenever the kernel $\kappa$ is universal, that is when the associated RKHS $\mathcal{H}$ is dense in the space of continuous functions on $\Omega$ [sriperumbudur2011universality]. One can show this condition is equivalent to an injectivity assumption on the operator $\Phi\star$.
Lemma A.1.
Let $\phi$ satisfy Assumption 3. Then $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ is injective if and only if $\mathcal{H}$ is dense in the space $\mathcal{C}^{0}(\Omega)$ of continuous functions over $\Omega$. In this case, $\mathrm{MMD}_{\kappa}$ is a distance on $\mathcal{M}(\Omega)$.
Proof.
The fact the $\mathrm{MMD}$ is a distance on $\mathcal{M}(\Omega)$ when $\Phi\star$ is injective directly follows from eq. 43. For the direct implication, assume $\Phi\star$ is injective and consider some measure $\nu\in\mathcal{H}^{\perp}$ i.e. such that for every $h\in\mathcal{H}$ we have $\int h\mathrm{d}\nu=0$. Then by the characterisation in eq. 41 we have for every $\alpha\in L^{2}(\rho)$:
$0=\int_{\Omega}\left(\int_{\mathbb{R}^{d}}\phi(\omega,x)\alpha(x)\mathrm{d}\rho(x)\right)\mathrm{d}\nu(\omega)=\langle\alpha,\Phi\star\nu\rangle_{L^{2}(\rho)}.$
Hence $\Phi\star\nu=0$, implying $\nu=0$ and thus that =$\mathcal{H}^{\perp}=\{0\}$ i.e. $\mathcal{H}$ is dense in $\mathcal{C}^{0}(\Omega)$. For the converse implication, assume that $\mathcal{H}$ is dense in $\mathcal{C}^{0}(\Omega)$ and consider some $\nu\in\mathcal{M}(\Omega)$ s.t. $\Phi\star\nu=0$. Then for every $\alpha\in L^{2}(\rho)$ we have $\langle\alpha,\Phi\star\nu\rangle_{L^{2}(\rho)}=0$ and by similar calculations this implies $\nu\in\mathcal{H}^{\perp}$ i.e. $\nu=0$. ∎
Kernel and integral operator
In this work we have used properties of the RKHS $\mathcal{H}$ seen as a subspace of the Hilbert space $L^{2}(\mu)$ for probability measures $\mu\in\mathcal{P}(\Omega)$. For such a probability measure $\mu\in\mathcal{P}(\Omega)$, it indeed follows from Assumption 3 that $\mathcal{H}$ is compactly embedded in $L^{2}(\mu)$ [steinwart2012mercer, Lem. 2.3]. Also, the kernel defines an integral operator $J_{\mu}:L^{2}(\mu)\to L^{2}(\mu)$ given by:
$\forall f\in L^{2}(\mu),\quad J_{\mu}\cdot f=\int_{\Omega}k(.,\omega)f(\omega)\mathrm{d}\mu(\omega).$
Then $J_{\mu}$ is compact, self-adjoint and positive operator and, by the spectral theorem, it can be diagonalized in an orthonormal basis $(e_{i})_{i\geq 0}$ of $L^{2}(\mu)$ with associated eigenvalues $(\lambda_{i})_{i\geq 0}$ s.t. $\lambda_{1}\geq\lambda_{2}\geq...\geq 0$. In particular, $J_{\mu}=\Phi_{\mu}^{*}\Phi_{\mu}$ with $\Phi_{\mu}:L^{2}(\mu)\to L^{2}(\rho)$ the feature operator defined in eq. 3, thus $(\sqrt{\lambda_{i}})_{i\geq 0}$ are the (right) singular values of $\Phi_{\mu}$ (which is a compact operator) and, if $\Phi\star$ is injective (i.e. $\mathcal{H}$ is universal according to Lemma A.1), then $\lambda_{i}>0$ for every $i\geq 0$. Mercer’s theorem gives a representation of the kernel $\kappa$ and of the associated RKHS $\mathcal{H}$ in terms of ths eigenvalue decomposition [steinwart2008support, Thm. 4.51].
Theorem A.1 (Mercer representation of RKHSs).
Assume $\Phi\star$ is injective and let $\mu\in\mathcal{P}(\Omega)$ be a probability measure with full support on $\Omega$. Consider $(\lambda_{i})_{i\geq 0}$ and $(e_{i})_{i\geq 0}$ the eigenvalue decomposition of the operator $J_{\mu}$. Then we have:
$\forall\omega,\omega^{\prime}\in\Omega\times\Omega,\quad\kappa(\omega,\omega^{\prime})=\sum_{i\geq 0}\lambda_{i}e_{i}(\omega)e_{i}(\omega^{\prime}),$
where the convergence is absolute and uniform over $\Omega\times\Omega$. Moreover
=$\mathcal{H}=\left\{\sum_{i\geq 0}a_{i}\sqrt{\lambda_{i}}e_{i},\,(a_{i})_{i\geq 0}\in\ell^{2}(\mathbb{N})\right\}$
is the RKHS associated to the kernel $\kappa$ when provided with scalar product $\langle.,.\rangle_{\mathcal{H}}$ defined for $f=\sum_{i\geq 0}a_{i}\sqrt{\lambda_{i}}$ and $g=\sum_{i\geq 0}b_{i}\sqrt{\lambda_{i}}$ by $\langle f,g\rangle_{\mathcal{H}}=\sum_{i\geq 0}a_{i}b_{i}$.